### v1

# Strong primeness for equivalence relations arising from Zariski dense subgroups

Abstract.
We show that orbit equivalence relations arising from essentially free ergodic probability measure preserving actions of Zariski dense discrete subgroups of simple algebraic groups are strongly prime. As a consequence, we prove the existence and the uniqueness of a prime factorization for orbit equivalence relations arising from direct products of higher rank lattices. This extends and strengthens Zimmer’s primeness result for equivalence relations arising from actions of lattices in simple Lie groups. The proof of our main result relies on a combination of ergodic theory of algebraic group actions and Popa’s intertwining theory for equivalence relations.
Key words and phrases:
Algebraic groups; Equivalence relations; Higher rank lattices; Strong primeness; von Neumann algebras; Zariski dense subgroups
2020 Mathematics Subject Classification:
37A20, 20G25, 46L10
DD was supported by Engineering and Physical Sciences Research Council grant EP/X026647/1
CH is supported by Institut Universitaire de France
1. Introduction and statement of the main results
Throughout, by an equivalence relation $\mathscr{R}$, we simply mean a probability measure preserving (pmp) countable Borel equivalence relation defined on a standard probability space $(X,\nu)$.
Introduction
We say that a type ${\rm II_{1}}$ ergodic equivalence relation $\mathscr{R}$ is prime if for any ergodic equivalence relations $\mathscr{T}_{1},\mathscr{T}_{2}$ for which $\mathscr{R}\cong\mathscr{T}_{1}\times\mathscr{T}_{2}$, then there exists $i\in\{1,2\}$ such that $\mathscr{T}_{i}$ has finite orbits almost everywhere. Zimmer’s seminal work [Zi81] shows that for any lattice $\Gamma<G$ in a noncompact simple connected real Lie group with finite center and any essentially free ergodic pmp action $\Gamma\curvearrowright(X,\nu)$, the orbit equivalence relation $\mathscr{R}(\Gamma\curvearrowright X)$ is prime.
In order to motivate and state our main results, we introduce the following terminology using the framework of Popa’s intertwining theory for equivalence relations (see Section 2 for further details). We say that a type ${\rm II_{1}}$ ergodic equivalence relation $\mathscr{R}$ is strongly prime if for any ergodic equivalence relations $\mathscr{S}$, $\mathscr{T}_{1}$, $\mathscr{T}_{2}$ for which ×$\mathscr{R}\times\mathscr{S}\cong\mathscr{T}_{1}\times\mathscr{T}_{2}$, then there exists $i\in\{1,2\}$ such that as subequivalence relations, $\mathscr{T}_{i}$ embeds into $\mathscr{S}$ inside ×$\mathscr{R}\times\mathscr{S}$, that we write $\mathscr{T}_{i}\preceq_{\mathscr{R}\times\mathscr{S}}\mathscr{S}$ (see Definition 2.4 for a precise statement). Note that this implies in particular that there exists an ergodic equivalence relation $\mathscr{U}$ such that $\mathscr{T}_{i}\times\mathscr{U}$ and $\mathscr{S}$ are stably isomorphic (see [Sp21]). We point out that the notion of strong primeness for type ${\rm II_{1}}$ factors was studied by Isono in [Is16].
The ergodic theory of equivalence relations is intimately connected to the theory of von Neumann algebras. By a well-known construction due to Feldman–Moore [FM75], to any ergodic equivalence relation $\mathscr{R}$ on a standard probability space $(X,\nu)$, one can associate a von Neumann factor $\operatorname{L}(\mathscr{R})$ so that $\operatorname{L}^{\infty}(X,\nu)\subset\operatorname{L}(\mathscr{R})$ sits as a Cartan subalgebra. Recall that a type ${\rm II_{1}}$ factor $M$ is prime if for any factors $M_{1},M_{2}$ for which $M\cong M_{1}\mathbin{\overline{\otimes}}M_{2}$, then there exists $i\in\{1,2\}$ such that $M_{i}$ is finite dimensional. For any ergodic equivalence relation $\mathscr{R}$, if $\operatorname{L}(\mathscr{R})$ is prime, then $\mathscr{R}$ is prime, but the converse need not hold in general (see Proposition 3.7 below). In relation with Zimmer’s primeness result [Zi81], in the case of simple connected real Lie groups with finite center and real rank one, any discrete subgroup $\Gamma<G$ is biexact in the sense of [BO08]. Then by Ozawa’s relative solidity result [Oz04], it follows that for any essentially free ergodic pmp action $\Gamma\curvearrowright(X,\nu)$ of a countable discrete nonamenable biexact group, the group measure space factor $\operatorname{L}(\Gamma\curvearrowright X)$ is prime. For other iconic primeness results in von Neumann algebra theory, we refer to [Ge96, Oz03, Pe06].
Over the last two decades, Popa’s deformation/rigidity theory (see the ICM surveys [Po06, Va10, Io18]) has led to a plethora of primeness and indecomposability results for von Neumann algebras and orbit equivalence relations arising from essentially free ergodic pmp actions of negatively curved groups such as free groups, hyperbolic groups, biexact groups, groups with a positive first $\ell^{2}$-betti number and so on (see also [Ad92, Ki05]). Except for very few results (see e.g. [DHI16, BIP18, IM19]), higher rank lattices remain allergic to deformation/rigidity theory methods.
The key novelty of our paper is to combine methods from ergodic theory of algebraic group actions and Popa’s intertwining theory to obtain strong primeness and existence and uniqueness of prime factorization for orbit equivalence relations arising from essentially free ergodic pmp actions of higher rank lattices and more generally Zariski dense discrete subgroups of simple algebraic groups.
Statement of the main results
Our first main theorem is the following strong primeness result strengthening Zimmer’s result [Zi81, Theorem 1.1].
Theorem A.
Let $k$ be a local field of characteristic zero, $\mathbf{G}$ a $k$-isotropic almost $k$-simple algebraic $k$-group and $\Gamma<\mathbf{G}(k)$ a Zariski dense discrete subgroup. Let $\Gamma\curvearrowright(X,\nu)$ be an essentially free ergodic pmp action.
Then the orbit equivalence relation $\mathscr{R}(\Gamma\curvearrowright X)$ is strongly prime.
By Borel’s density theorem [Bo60], Theorem A applies to all essentially free ergodic pmp actions of lattices $\Gamma<\mathbf{G}(k)$, where $\mathbf{G}$ is a $k$-isotropic almost $k$-simple algebraic $k$-group (e.g. $\operatorname{SL}_{n}(\mathbb{Z})<\operatorname{SL}_{n}(\mathbb{R})$ for every $n\geq 2$).
In that respect, denote by $\mathfrak{Z}$ the class of orbit equivalence relations $\mathscr{R}=\mathscr{R}(\Gamma\curvearrowright X)$ that arise from essentially free ergodic pmp actions $\Gamma\curvearrowright(X,\nu)$, where $\Gamma<\mathbf{G}(k)$ is a Zariski dense discrete subgroup, $k$ is a local field of characteristic zero and $\mathbf{G}$ is a $k$-isotropic almost $k$-simple algebraic $k$-group.
Using Theorem A, we deduce the following unique prime factorization result for direct products of ergodic equivalence relations that belong to $\mathfrak{Z}$.
Corollary B.
Let $n\geq 2$ and $\mathscr{R}_{1},\dots,\mathscr{R}_{n}$ be ergodic equivalence relations that belong to $\mathfrak{Z}$. Then the following assertions hold:
  • $(\rm i)$
    If $\mathscr{R}=\mathscr{R}_{1}\times\cdots\times\mathscr{R}_{n}\cong\mathscr{T}_{1}\times\mathscr{T}_{2}$, where $\mathscr{T}_{1},\mathscr{T}_{2}$ are type ${\rm II_{1}}$ ergodic equivalence relations, then there exists a partition $\{1,\dots,n\}=T_{1}\sqcup T_{2}$ into nonempty subsets such that for every $j\in\{1,2\}$, we have $\mathscr{T}_{j}\preceq_{\mathscr{R}}\mathscr{R}_{T_{j}}$ as subequivalence relations, where $\mathscr{R}_{T_{j}}=\prod_{i\neq j}\mathscr{R}_{i}$.
  • $(\rm ii)$
    If $\mathscr{R}=\mathscr{R}_{1}\times\cdots\times\mathscr{R}_{n}\cong\mathscr{T}_{1}\times\cdots\times\mathscr{T}_{p}$, where $\mathscr{T}_{1},\dots,\mathscr{T}_{p}$ are type ${\rm II_{1}}$ ergodic equivalence relations and $p\geq n$, then $n=p$ and upon permuting the indices, $\mathscr{T}_{i}$ and $\mathscr{R}_{i}$ are stably isomorphic for every $1\leq i\leq n$.
We point out that the first unique prime factorization result appeared in von Neumann algebra theory in the pioneering work of Ozawa–Popa [OP03]. Indeed, they showed that tensor products of type ${\rm II_{1}}$ factors arising from countable discrete icc nonamenable biexact groups have a unique tensor product decomposition upon taking amplifications and permuting the indices.
Assuming that the equivalence relations $\mathscr{R}_{1},\dots,\mathscr{R}_{n}$ that belong to $\mathfrak{Z}$ are moreover strongly ergodic, we obtain the following sharper unique prime factorization result.
Corollary C.
Let $n\geq 2$ and $\mathscr{R}_{1},\dots,\mathscr{R}_{n}$ be strongly ergodic equivalence relations that belong to $\mathfrak{Z}$. Then the following assertions hold:
  • $(\rm i)$
    If $\mathscr{R}=\mathscr{R}_{1}\times\cdots\times\mathscr{R}_{n}\cong\mathscr{T}_{1}\times\mathscr{T}_{2}$, where $\mathscr{T}_{1},\mathscr{T}_{2}$ are type ${\rm II_{1}}$ ergodic equivalence relations, then there exist a partition $\{1,\dots,n\}=T_{1}\sqcup T_{2}$ into nonempty subsets and $t>0$ such that
    $\mathscr{T}_{1}^{t}\cong\mathscr{R}_{T_{1}}\quad\text{and}\quad\mathscr{T}_{2}^{1/t}\cong\mathscr{R}_{T_{2}}.$
  • $(\rm ii)$
    If $\mathscr{R}=\mathscr{R}_{1}\times\cdots\times\mathscr{R}_{n}\cong\mathscr{T}_{1}\times\cdots\times\mathscr{T}_{p}$, where $\mathscr{T}_{1},\dots,\mathscr{T}_{p}$ are type ${\rm II_{1}}$ ergodic equivalence relations and $p\geq n$, then $n=p$ and there exist $t_{1},\dots,t_{n}>0$ such that $t_{1}\cdots t_{n}=1$ and upon permuting the indices, $\mathscr{T}_{i}^{t_{i}}\cong\mathscr{R}_{i}$ for every $1\leq i\leq n$.
In particular, by Borel’s density theorem [Bo60] and Kazhdan’s theorem [Ka66], Corollary C applies to direct products of equivalence relations arising from essentially free ergodic pmp actions of lattices $\Gamma<\mathbf{G}(k)$, where $\mathbf{G}$ is an almost $k$-simple algebraic $k$-group such that $\operatorname{rk}_{k}(\mathbf{G})\geq 2$ (e.g. $\operatorname{SL}_{n}(\mathbb{Z})<\operatorname{SL}_{n}(\mathbb{R})$ for every $n\geq 3$). Corollary C should be compared with Hoff’s unique prime factorization result for strongly ergodic equivalence relations with nontrivial $1$-cohomology [Ho15]. Let us mention that Isono–Marrakchi [IM19] obtained a unique prime factorization result for tensor products of group measure space factors arising from compact pmp actions of higher rank lattices with Kazhdan’s property (T). In that respect, Corollaries B and C provide the first unique prime factorization result for orbit equivalence relations arising from arbitrary pmp actions of higher rank lattices.
Our second main theorem gives a complete characterization of primeness for orbit equivalence relations arising from essentially free ergodic pmp actions of product groups $\Gamma_{1}\times\cdots\times\Gamma_{n}\curvearrowright(X,\nu)$, where $\Gamma_{i}$ is a discrete group as in Theorem A with Kazhdan’s property (T).
Theorem D.
Let $n\geq 1$. For every $1\leq i\leq n$, let $k_{i}$ be a local field of characteristic zero, $\mathbf{G}_{i}$ a $k_{i}$-isotropic $k_{i}$-simple algebraic $k_{i}$-group and $\Gamma_{i}<\mathbf{G}_{i}(k_{i})$ a Zariski dense discrete subgroup with Kazhdan’s property (T). Set $\Gamma=\Gamma_{1}\times\cdots\times\Gamma_{n}$. Let $\Gamma\curvearrowright(X,\nu)$ be an essentially free ergodic pmp action and set $\mathscr{R}=\mathscr{R}(\Gamma\curvearrowright X)$. Assume that $\mathscr{R}\cong\mathscr{T}_{1}\times\mathscr{T}_{2}$, where $\mathscr{T}_{1},\mathscr{T}_{2}$ are type ${\rm II_{1}}$ ergodic equivalence relations.
Then there exist a partition $\{1,\dots,n\}=T_{1}\sqcup T_{2}$ into nonempty subsets, $t>0$, an essentially free ergodic pmp action $\Gamma_{T_{j}}\curvearrowright(X_{j},\nu_{j})$, where $\Gamma_{T_{j}}=\prod_{i\in T_{j}}\Gamma_{i}$ for every $j\in\{1,2\}$, such that the following assertions hold:
  • $(\rm i)$
    $\Gamma\curvearrowright(X,\nu)\cong\Gamma_{T_{1}}\times\Gamma_{T_{2}}\curvearrowright(X_{1}\times X_{2},\nu_{1}\otimes\nu_{2})$ and
  • $(\rm ii)$
    $\mathscr{T}_{1}^{t}\cong\mathscr{R}(\Gamma_{T_{1}}\curvearrowright X_{1})$ and $\mathscr{T}_{2}^{1/t}\cong\mathscr{R}(\Gamma_{T_{2}}\curvearrowright X_{2})$.
Theorem D is an analogue of Drimbe’s primeness characterization for group measure space factors arising from essentially free ergodic pmp actions of products of hyperbolic groups with Kazhdan’s property (T) [Dr19]. By Borel’s density theorem [Bo60] and Kazhdan’s theorem [Ka66], Theorem D applies to all essentially free ergodic pmp actions $\Gamma\curvearrowright(X,\nu)$, where $\Gamma_{i}<\mathbf{G}_{i}(k_{i})$ is a lattice and $\mathbf{G}_{i}$ is a $k_{i}$-simple algebraic $k_{i}$-group such that $\operatorname{rk}_{k_{i}}(\mathbf{G}_{i})\geq 2$ for every $i\in\{1,\dots,n\}$.
As a consequence of Theorem D, we obtain the following existence and uniqueness of a prime factorization for equivalence relations.
Corollary E.
Let $n\geq 1$. For every $1\leq i\leq n$, let $k_{i}$ be a local field of characteristic zero, $\mathbf{G}_{i}$ a $k_{i}$-isotropic $k_{i}$-simple algebraic $k_{i}$-group and $\Gamma_{i}<\mathbf{G}_{i}(k_{i})$ a Zariski dense discrete subgroup with Kazhdan’s property (T). Set $\Gamma=\Gamma_{1}\times\cdots\times\Gamma_{n}$. Let $\Gamma\curvearrowright(X,\nu)$ be an essentially free ergodic pmp action.
Then there exist a unique partition $\{1,\dots,n\}=T_{1}\sqcup\cdots\sqcup T_{r}$ into nonempty subsets (upon permuting the indices), an essentially free ergodic pmp action $\Gamma_{T_{j}}\curvearrowright(X_{j},\nu_{j})$, where $\Gamma_{T_{j}}=\prod_{i\in T_{j}}\Gamma_{i}$ for every $j\in\{1,\dots,r\}$, such that the following assertions hold:
  • $(\rm i)$
    $\Gamma\curvearrowright(X,\nu)\cong\Gamma_{T_{1}}\times\cdots\times\Gamma_{T_{r}}\curvearrowright(X_{1}\times\cdots\times X_{r},\nu_{1}\otimes\cdots\otimes\nu_{r})$ and
  • $(\rm ii)$
    $\mathscr{R}(\Gamma_{T_{j}}\curvearrowright X_{j})$ is prime for every $j\in\{1,\dots,r\}$.
Comments on the proofs
The proof of Theorem A builds upon Zimmer’s proof of [Zi81, Theorem 1.1] and combines ergodic theory of algebraic group actions and Popa’s intertwining theory for equivalence relations. Unlike Zimmer’s proof of [Zi81, Theorem 1.1], we do not use induction and we work directly with the orbit equivalence relation $\mathscr{R}(\Gamma\curvearrowright X)$. This allows us to deal with essentially free ergodic pmp action of arbitrary Zariski dense discrete subgroups $\Gamma<\mathbf{G}(k)$ rather than lattices $\Gamma<\mathbf{G}(k)$ as in [Zi81]. The proof of Corollary B follows by combining Theorem A with Popa’s intertwining theory.
Using Tucker-Drob’s results on the structure of inner amenable groups [TD14], we observe in Proposition 2.7 that any Zariski dense discrete subgroup $\Gamma<\mathbf{G}(k)$ as in Theorem A is not inner amenable. This implies that for any essentially free strongly ergodic pmp action $\Gamma\curvearrowright(X,\nu)$, the group measure space von Neumann factor $\operatorname{L}(\Gamma\curvearrowright X)$ is full meaning that it has no nontrivial central sequences (see [Ch81]). For the proof of Corollary C, we use Theorem A and we exploit Popa’s intertwining theory in the setting of von Neumann algebras in combination with Isono–Marrakchi’s results on tensor product decompositions of full factors [IM19].
The proof of Theorem D and Corollary E follows the proof of Theorem A and moreover exploits intertwining techniques from [Dr19].
Acknowledgments
This work was initiated when CH was visiting the Mathematical Institute of the University of Oxford during Hilary Term 2024. He is grateful towards Stuart White for his kind invitation. We also thank Adrian Ioana for his valuable comments.
2. Preliminaries
2.1. Equivalence relations
Let $\mathscr{R}$ be an equivalence relation defined on a standard probability space $(X,\nu)$. Denote by $\mathscr{B}(X)$ (resp. $\mathscr{B}(\mathscr{R})$) the $\sigma$-algebra of all Borel subsets of $X$ (resp. $\mathscr{R}$). For every $x\in X$, we denote by $[x]_{\mathscr{R}}$ the $\mathscr{R}$-equivalence class of $x\in X$. We define the $\sigma$-finite Borel measure $m$ on $\mathscr{R}$ by the formula
$\forall\mathscr{W}\in\mathscr{B}(\mathscr{R}),\quad m(\mathscr{W})=\int_{X}\sharp(\left\{(x,y)\mid y\in[x]_{\mathscr{R}}\right\}\cap\mathscr{W})\,{\rm d}\nu(x).$
We denote by $\operatorname{Aut}(X,\nu)$ the group of all pmp Borel automorphisms $\theta:(X,\nu)\to(X,\nu)$. We denote by $\operatorname{Aut}(\mathscr{R})$ the automorphism group of $\mathscr{R}$ that consists of all pmp Borel automorphisms $\theta\in\operatorname{Aut}(X,\nu)$ such that $(\theta(x),\theta(y))\in\mathscr{R}$ if and only if $(x,y)\in\mathscr{R}$ for $(x,y)\in X\times X$. Then we denote by $[\mathscr{R}]<\operatorname{Aut}(\mathscr{R})$ the full group of $\mathscr{R}$ that consists of all automorphisms $\theta\in\operatorname{Aut}(\mathscr{R})$ for which $(\theta(x),x)\in\mathscr{R}$ for $\nu$-almost every $x\in X$. The uniform metric
$d_{u}:[\mathscr{R}]\times[\mathscr{R}]\to\mathbb{R}_{+}:(\theta,\rho)\mapsto\nu(\{x\in X\mid\theta(x)\neq\rho(x)\})$
is complete and separable and so the full group $[\mathscr{R}]$ is a Polish group. By [FM75], there exists a countable subgroup $\Lambda<[\mathscr{R}]$ such that $\mathscr{R}=\mathscr{R}(\Lambda\curvearrowright X)$ is the orbit equivalence relation of the pmp action $\Lambda\curvearrowright(X,\nu)$. The full pseudogroup of $\mathscr{R}$ consists of all pmp Borel partial isomorphisms $\theta:U\to V$, where $U,V\in\mathscr{B}(X)$ and $(\theta(x),x)\in\mathscr{R}$ for $\nu$-almost every $x\in U$. For every $\theta:U\to V\in[[\mathscr{R}]]$, there exists $\rho\in[\mathscr{R}]$ such that $\rho|_{U}=\theta$. We say that $\mathscr{R}$ is ergodic if and only if the Koopman representation $\kappa:[\mathscr{R}]\to\mathscr{U}(\operatorname{L}^{2}(X,\nu)^{0})$ is ergodic, where $\operatorname{L}^{2}(X,\nu)^{0}=\operatorname{L}^{2}(X,\nu)\ominus\mathbb{C}\mathbf{1}_{X}$.
Whenever $U\subset X$ is a nonnull measurable subset, we define the equivalence relation $\mathscr{R}|_{U}=\mathscr{R}\cap(U\times U)$ on the standard probability space $(U,\nu_{U})$ where $\nu_{U}=\frac{1}{\nu(U)}\nu|_{U}$. Assume that $\mathscr{R}$ is ergodic. For every $t>0$, we define the amplification $\mathscr{R}^{t}$ as follows. Denote by $c$ the counting measure on $\mathbb{N}$. Choose a measurable subset $X_{t}\subset X\times\mathbb{N}$ such that $(\nu\otimes c)(X_{t})=t$ and set $\nu_{t}=\frac{1}{(\nu\otimes c)(X_{t})}(\nu\otimes c)|_{X_{t}}$. Define the ergodic equivalence relation $\mathscr{R}^{t}$ on the standard probability space $(X_{t},\nu_{t})$ by declaring $((x,p),(y,q))\in\mathscr{R}^{t}$ if and only if $(x,y)\in\mathscr{R}$ for $((x,p),(y,q))\in X_{t}\times X_{t}$.
For every $n\in\mathbb{N}\cup\{+\infty\}$, define the Borel subset
$X_{n}=\left\{x\in X\mid\sharp([x]_{\mathscr{R}})=n\right\}\subset X.$
We say that $\mathscr{R}$ is (essentially) finite if $\nu(X_{\infty})=0$. We say that $\mathscr{R}$ is (essentially) bounded if there exists $k\in\mathbb{N}$ such that $\nu(X_{n})=0$ for every $n\geq k$.
Let $\mathscr{S}$ be another equivalence relation defined on a standard probability space $(Y,\eta)$. We say that $\mathscr{R}$ and $\mathscr{S}$ are isomorphic if there exists a pmp Borel isomorphism $\theta:(X,\nu)\to(Y,\eta)$ such that for $m$-almost every $(x,y)\in\mathscr{R}$ we have $(\theta(x),\theta(y))\in\mathscr{S}$. We say that $\mathscr{R}$ and $\mathscr{S}$ are stably isomorphic if there exist nonnull measurable subsets $U\subset X$ and $V\subset Y$ and a pmp Borel isomorphism $\theta:(U,\nu_{U})\to(V,\eta_{V})$ such that for $m$-almost every $(x,y)\in\mathscr{R}|_{U}$, we have $(\theta(x),\theta(y))\in\mathscr{S}|_{V}$.
Let $\mathscr{S}\leq\mathscr{R}$ be a subequivalence relation. For every $n\in\mathbb{N}\cup\{+\infty\}$, define the Borel subset
$Y_{n}=\left\{x\in X\mid[x]_{\mathscr{R}}\text{ is the union of }n\;\mathscr{S}\text{-classes}\right\}\subset X.$
We say that $\mathscr{S}\leq\mathscr{R}$ has (essentially) finite index if $\nu(Y_{\infty})=0$. We say that $\mathscr{S}\leq\mathscr{R}$ has (essentially) bounded index if there exists $k\in\mathbb{N}$ such that $\nu(Y_{n})=0$ for every $n\geq k$.
Following [FM75], we denote by $\operatorname{L}(\mathscr{R})$ the tracial von Neumann algebra associated with $\mathscr{R}$. Then $\operatorname{L}^{\infty}(X)\subset\operatorname{L}(\mathscr{R})$ is a Cartan subalgebra. We simply denote by $\left\{u_{\theta}\mid\theta\in[\mathscr{R}]\right\}\subset\mathscr{U}(\operatorname{L}(\mathscr{R}))$ the full group of $\mathscr{R}$ regarded as a subgroup of the unitary group $\mathscr{U}(\operatorname{L}(\mathscr{R}))$. Then we have
$\operatorname{L}(\mathscr{R})=\left\{u_{\theta}\mid\theta\in[\mathscr{R}]\right\}^{\prime\prime}\vee\operatorname{L}^{\infty}(X).$
Facts 2.1.
Keep the same notation as above. We will be using the following useful properties.
  • $(\rm i)$
    The equivalence relation $\mathscr{R}$ is ergodic if and only if the Koopman representation $\kappa:[\mathscr{R}]\to\mathscr{U}(\operatorname{L}^{2}(X,\nu)^{0})$ is weakly mixing, where $\operatorname{L}^{2}(X,\nu)^{0}=\operatorname{L}^{2}(X,\nu)\ominus\mathbb{C}\mathbf{1}_{X}$.
  • $(\rm ii)$
    If $\mathscr{R}$ is ergodic and $(X,\nu)$ is diffuse, then $\operatorname{L}(\mathscr{R})=\left\{u_{\theta}\mid\theta\in[\mathscr{R}]\right\}^{\prime\prime}$.
  • $(\rm iii)$
    Let $\Psi\in\operatorname{Aut}(X,\nu)$ be a pmp Borel automorphism and denote by $\mathscr{R}_{\Psi}$ the equivalence relation generated by $\Psi$. Then for any $\varepsilon>0$, there exist a measurable subset $X_{0}\subset X$ with $\nu(X\setminus X_{0})\leq\varepsilon$, a positive integer $p$ and $\Psi_{0}\in[\mathscr{R}_{\Psi}]$ such that $\Psi_{0}=\Psi$ on $X_{0}$ and $\Psi_{0}^{p}=\operatorname{id}_{X}$.
Proof.
$(\rm i)$ This follows from [LM24, Theorem 11.20].
$(\rm ii)$ Indeed, it suffices to prove that $\operatorname{L}^{\infty}(X)\subset\left\{u_{\theta}\mid\theta\in[\mathscr{R}]\right\}^{\prime\prime}$. Let $Y\subset X$ be a measurable subset such that $\nu(Y)<1$. Set $Z=X\setminus Y$ and $\mathscr{S}=\mathscr{R}|_{Z}$. Since $\mathscr{S}$ is ergodic and $(Z,\nu_{Z})$ is diffuse, there exists an ergodic type ${\rm II_{1}}$ hyperfinite subequivalence relation $\mathscr{T}\leq\mathscr{S}$ (see e.g. [Zi84, Proposition 9.3.2]). By [CFW81], choose a free ergodic pmp action $\mathbb{Z}\curvearrowright(Z,\nu_{Z})$ such that $\mathscr{T}=\mathscr{R}(\mathbb{Z}\curvearrowright Z)$. Denote by $\rho\in[\mathscr{T}]$ the generator of $\mathbb{Z}\curvearrowright(Z,\nu_{Z})$. For every $n\in\mathbb{N}$, define $\theta_{n}\in[\mathscr{R}]$ by $\theta_{n}(x)=x$ if $x\in Y$ and $\theta_{n}(x)=\rho^{n}(x)$ if $x\in Z$. Then it is plain to see that $u_{\theta_{n}}\to\mathbf{1}_{Y}$ weakly as $n\to\infty$. This shows that $\operatorname{L}(\mathscr{R})=\left\{u_{\theta}\mid\theta\in[\mathscr{R}]\right\}^{\prime\prime}$.
$(\rm iii)$ Since $\mathscr{R}_{\Psi}$ is hyperfinite by Rokhlin’s lemma, it follows that there exists an increasing sequence of finite equivalence relations $\mathscr{T}_{1}\leq\mathscr{T}_{2}\leq\dots$ such that (up to a conull Borel subset) we have $\mathscr{R}_{\Psi}=\bigcup_{k\geq 1}\mathscr{T}_{k}$. Then for $\nu$-almost every $x\in X$, we have $[x]_{\mathscr{R}_{\Psi}}=\bigcup_{k\geq 1}[x]_{\mathscr{T}_{k}}$, and thus, there exists $k\geq 1$ such that $\Psi(x)\in[x]_{\mathscr{T}_{k}}$. Next, define the measurable sets $X_{k}=\{x\in X\mid\Psi(x)\in[x]_{\mathscr{T}_{k}}\}$, $k\geq 1$ and note that $X=\bigcup_{k\geq 1}X_{k}$. Take $N\geq 1$ such that $\nu(X\setminus X_{N})\leq\varepsilon$ and let $\Psi_{0}\in[\mathscr{T}_{N}]\leq[\mathscr{R}_{\Psi}]$ be such that $\Psi_{0}=\Psi$ on $X_{N}$. By construction, it follows that the equivalence relation generated by $\Psi_{0}$ is finite, hence, the lemma follows by letting $X_{0}=X_{N}$. ∎
2.2. Popa’s intertwining theory for equivalence relations
We review Popa’s criterion for intertwining von Neumann subalgebras [Po01, Po03]. Let $(M,\tau)$ be a tracial von Neumann algebra and $A\subset 1_{A}M1_{A}$, $B\subset 1_{B}M1_{B}$ be von Neumann subalgebras. Let $\mathcal{G}\subset\mathscr{U}(A)$ be a subgroup such that $\mathscr{G}^{\prime\prime}=A.$ By [Po03, Corollary 2.3] and [Po01, Theorem A.1] (see also [Va06, Proposition C.1]), the following conditions are equivalent:
  • $(\rm i)$
    There exist $d\geq 1$, a projection $q\in\operatorname{M}_{d}(B)$, a nonzero partial isometry $v\in\operatorname{M}_{1,d}(1_{A}M)q$ and a unital normal $\ast$-homomorphism $\pi:A\to q\operatorname{M}_{d}(B)q$ such that $av=v\pi(a)$ for all $a\in A$.
  • $(\rm ii)$
    There is no net of unitaries $(w_{n})_{n}$ in $\mathscr{G}$ such that
    $\forall x,y\in 1_{A}M1_{B},\quad\lim_{n}\|\operatorname{E}_{B}(x^{*}w_{n}y)\|_{2}=0.$
If one of the previous equivalent conditions is satisfied, we say that $A$ embeds into $B$ inside $M$ and write $A\preceq_{M}B$. Moreover, if the inclusion $B\subset M$ is unital, then the above are also equivalent to:
  • $(\rm iii)$
    The unitary representation $\pi:\mathscr{G}\to\mathscr{U}(\operatorname{L}^{2}(\langle M,e_{B}\rangle,\operatorname{Tr}))$ given by $\pi_{u}(\xi)=u\xi u^{*}$ for $u\in\mathscr{G}$, $\xi\in\operatorname{L}^{2}(\langle M,e_{B}\rangle,\operatorname{Tr})$, is not weakly mixing. Here, we denoted by ${\rm Tr}$ the associated semifinite trace on Jones’ basic construction $\langle M,e_{B}\rangle$.
If for every nonzero projection $p\in A^{\prime}\cap 1_{A}M1_{A}$ we have $Ap\preceq_{M}B$, then we write $A\preceq^{s}_{M}B$.
Popa’s criterion for intertwining von Neumann subalgebras was adapted to the setting of equivalence relations by Ioana [Io11]. Let $\mathscr{R}$ be an equivalence relation defined on a standard probability space $(X,\nu)$. Let $Y,Z\subset X$ be nonnull measurable subsets and $\mathscr{S}\leq\mathscr{R}|_{Y}$ and $\mathscr{T}\leq\mathscr{R}|_{Z}$ be subequivalence relations. Following [IKT08], define the map
$\varphi_{\mathscr{T}}:[[\mathscr{R}]]\to[0,1]:\theta\mapsto\nu(\{x\in\operatorname{dom}(\theta)\mid(\theta(x),x)\in\mathscr{T}\}).$
By [Io11, Lemma 1.7] (see also [Sp21, Lemma 3.1]), the following conditions are equivalent:
  • $(\rm i)$
    There exist a $\mathscr{S}$-invariant nonnull measurable subset $Y_{0}\subset Y$ and a subequivalence relation $\mathscr{S}_{0}\leq\mathscr{S}$ such that for any nonnull measurable subset $Y_{1}\subset Y_{0}$, there is a nonnull measurable subset $U\subset Y_{1}$ and $\theta\in[[\mathscr{R}]]$ with $\theta:U\to V$, such that
    • $\mathscr{S}_{0}|_{U}\leq\mathscr{S}|_{U}$ has bounded index, and
    • $(\theta\times\theta)(\mathscr{S}_{0}|_{U})\leq\mathscr{T}|_{V}$.
  • $(\rm ii)$
    There is no sequence $(\theta_{n})_{n}$ in $[\mathscr{S}]$ such that
    $\forall\psi,\rho\in[[\mathscr{R}]],\quad\lim_{n}\varphi_{\mathscr{T}}(\psi\theta_{n}\rho)=0.$
If one of the previous equivalent conditions is satisfied, we say that $\mathscr{S}$ embeds into $\mathscr{T}$ inside $\mathscr{R}$ and write $\mathscr{S}\preceq_{\mathscr{R}}\mathscr{T}$.
Facts 2.2.
Keep the same notation as above. We will be using the following useful properties.
  • $(\rm i)$
    We have $\mathscr{S}\preceq_{\mathscr{R}}\mathscr{T}$ if and only $\operatorname{L}(\mathscr{S})\preceq_{\operatorname{L}(\mathscr{R})}\operatorname{L}(\mathscr{T})$.
  • $(\rm ii)$
    If $\mathscr{S}_{0}\leq\mathscr{S}$ is a subequivalence relation and $\mathscr{S}\preceq_{\mathscr{R}}\mathscr{T}$, then we have $\mathscr{S}_{0}\preceq_{\mathscr{R}}\mathscr{T}$.
For every $i\in\{1,2\}$, let $(Z_{i},\nu_{i})$ be a standard probability space such that $(X,\nu)\cong(Z_{1}\times Z_{2},\nu_{1}\otimes\nu_{2})$. Let $Z\subset Z_{1}$ be a nonnull measurable subset and consider $\nu_{Z}=\frac{1}{\nu_{1}(Z)}\nu_{1}|_{Z}$. Let $\mathscr{S}$ be an ergodic equivalence relation on $(Z,\nu_{Z})$ such that $\mathscr{S}\times\operatorname{id}_{Z_{2}}\leq\mathscr{R}|_{Z\times Z_{2}}$. Denote by $\kappa:[\mathscr{S}]\to\mathscr{U}(\operatorname{L}^{2}(Z,\nu_{Z})^{0})$ the Koopman representation. Let $\mathscr{T}\leq\mathscr{R}$ be a subequivalence relation.
  • $(\rm iii)$
    Assume that $\mathscr{S}\times\operatorname{id}_{Z_{2}}\npreceq_{\mathscr{R}}\mathscr{T}$. Then for all $\varepsilon>0$, $n\geq 1$, $f_{1},\dots,f_{n}\in\operatorname{L}^{2}(Z,\nu_{Z})^{0}$ and $\psi_{1},\dots,\psi_{n}\in[\mathscr{R}]$, there exists $\theta\in[\mathscr{S}]$ such that $|\langle\kappa_{\theta}(f_{i}),f_{j}\rangle|\leq\varepsilon$ and $\varphi_{\mathscr{T}}(\psi_{i}(\theta\times\operatorname{id}_{Z_{2}})\psi_{j})\leq\varepsilon$ for all $1\leq i,j\leq n$.
For every $i\in\{1,2\}$, let $\mathscr{R}_{i}$ be an ergodic equivalence relation on $(X_{i},\nu_{i})$ and $\mathscr{S}_{i}$ an ergodic equivalence relation on $(Y_{i},\eta_{i})$. Assume that $\mathscr{R}_{1}\times\mathscr{S}_{1}\cong\mathscr{R}_{2}\times\mathscr{S}_{2}$. Set $(Z,\zeta)=(X_{1}\times Y_{1},\nu_{1}\otimes\eta_{1})\cong(X_{2}\times Y_{2},\nu_{2}\otimes\eta_{2})$ and $\mathscr{R}=\mathscr{R}_{1}\times\mathscr{S}_{1}\cong\mathscr{R}_{2}\times\mathscr{S}_{2}$. We simply write $\mathscr{R}_{1}\preceq_{\mathscr{R}}\mathscr{R}_{2}$ as subequivalence relations if $\mathscr{R}_{1}\times\operatorname{id}_{Y_{1}}\preceq_{\mathscr{R}}\mathscr{R}_{2}\times\operatorname{id}_{Y_{2}}$.
  • $(\rm iv)$
    If $\mathscr{S}_{1}\preceq_{\mathscr{R}}\mathscr{S}_{2}$ as subequivalence relations, then we have
    $\operatorname{L}^{\infty}(X_{2})\preceq_{\operatorname{L}(\mathscr{R})}\operatorname{L}^{\infty}(X_{1}).$
  • $(\rm v)$
    If $\mathscr{T}_{1}\leq\mathscr{S}_{1}$ is an ergodic subequivalence relation and $\mathscr{T}_{1}\preceq_{\mathscr{R}}\mathscr{S}_{2}$ as subequivalence relations, then we have $\mathscr{S}_{1}\preceq_{\mathscr{R}}\mathscr{S}_{2}$ as subequivalence relations.
  • $(\rm vi)$
    If $\mathscr{R}_{1}\preceq_{\mathscr{R}}\mathscr{R}_{2}$ as subequivalence relations, then there exists an ergodic equivalence relation $\mathscr{V}$ such that $\mathscr{R}_{1}\times\mathscr{V}$ and $\mathscr{R}_{2}$ are stably isomorphic.
Proof.
$(\rm i)$ This follows from [Io11, Lemma 1.8].
$(\rm ii)$ It is obvious from the definition.
$(\rm iii)$ Since $\mathscr{S}$ is ergodic, it follows by Facts 2.1(i),$(\rm i),(\rm ii)$ that $\operatorname{L}(\mathscr{S})=[\mathscr{S}]^{\prime\prime}$ and that the Koopman representation $\kappa:[\mathscr{S}]\to\mathscr{U}(\operatorname{L}^{2}(Z,\nu_{Z})^{0})$ is weakly mixing. Denote by $\left\{u_{\theta}\mid\theta\in[\mathscr{R}]\right\}\subset\mathscr{U}(\operatorname{L}(\mathscr{R}))$ the full group of $\mathscr{R}$ regarded as a subgroup of the unitary group $\mathscr{U}(\operatorname{L}(\mathscr{R}))$. By assumption, we have $\operatorname{L}(\mathscr{S})\mathbin{\overline{\otimes}}\operatorname{L}^{\infty}(Z_{2})\npreceq_{\operatorname{L}(\mathscr{R})}\operatorname{L}(\mathscr{T})$, which implies that there exist two sequences of unitaries $(v_{n})_{n\geq 1}\subset\mathscr{U}(\operatorname{L}(\mathscr{S}))$ and $(w_{n})_{n\geq 1}\subset\mathscr{U}(\operatorname{L}^{\infty}(Z_{2}))$ such that $\|\operatorname{E}_{\operatorname{L}(\mathscr{T})}(xv_{n}w_{n}u_{\theta})\|_{2}\to 0$, for all $x\in\operatorname{L}(\mathscr{R})$ and $\theta\in[\mathscr{R}]$. Since $u_{\theta}^{*}w_{n}u_{\theta}=w_{n}\circ\theta\in\mathscr{U}(\operatorname{L}^{\infty}(X))$, for any $n\geq 1$, and $\operatorname{L}^{\infty}(X)\subset\operatorname{L}(\mathscr{T})$, it follows that $\|\operatorname{E}_{\operatorname{L}(\mathscr{T})}(xv_{n}au_{\theta})\|_{2}\to 0$, for all $x\in\operatorname{L}(\mathscr{R})$, $a\in\operatorname{L}^{\infty}(X)$ and $\theta\in[\mathscr{R}]$. This implies that $\operatorname{L}(\mathscr{S})\npreceq_{\operatorname{L}(\mathscr{R})}\operatorname{L}(\mathscr{T})$. This further implies that the unitary representation $\pi:[\mathscr{S}]\to\mathscr{U}(\operatorname{L}^{2}(\langle\operatorname{L}(\mathscr{R}),e_{\operatorname{L}(\mathscr{T})}\rangle,\operatorname{Tr}))$ given by $\pi_{u}(\xi)=u\xi u^{*}$ for $u\in[\mathscr{S}]$, $\xi\in\operatorname{L}^{2}(\langle\operatorname{L}(\mathscr{R}),e_{\operatorname{L}(\mathscr{T})}\rangle,\operatorname{Tr})$, is weakly mixing. Since $\|\operatorname{E}_{\operatorname{L}(\mathscr{T})}(u_{\theta})\|_{2}^{2}=\varphi_{\mathscr{T}}(\theta)$ for any $\theta\in[\mathscr{R}]$ and since the direct sum $\kappa\oplus\pi$ is a weakly mixing unitary representation of $[\mathscr{S}]$ as well, the conclusion follows.
$(\rm iv)$ This follows from [Sp21, Corollary 3.4]
$(\rm v)$ If $\mathscr{T}_{1}\leq\mathscr{S}_{1}$ is an ergodic subequivalence relation and $\mathscr{T}_{1}\preceq_{\mathscr{R}}\mathscr{S}_{2}$ as subequivalence relations, then we have $\operatorname{L}^{\infty}(X_{1})\mathbin{\overline{\otimes}}\operatorname{L}(\mathscr{T}_{1})\preceq_{\operatorname{L}(\mathscr{R})}\operatorname{L}^{\infty}(X_{2})\mathbin{\overline{\otimes}}\operatorname{L}(\mathscr{S}_{2})$ and so $\operatorname{L}^{\infty}(X_{2})\preceq_{\operatorname{L}(\mathscr{R})}\operatorname{L}^{\infty}(X_{1})$ by [Va07, Lemma 3.5]. By applying again [Va07, Lemma 3.5], we infer that $\operatorname{L}^{\infty}(X_{1})\mathbin{\overline{\otimes}}\operatorname{L}(\mathscr{S}_{1})\preceq_{\operatorname{L}(\mathscr{R})}\operatorname{L}^{\infty}(X_{2})\mathbin{\overline{\otimes}}\operatorname{L}(\mathscr{S}_{2})$ and so $\mathscr{S}_{1}\preceq_{\mathscr{R}}\mathscr{S}_{2}$ as subequivalence relations.
$(\rm vi)$ This follows from [Sp21, Proposition 3.6]. ∎
The following result and its proof are inspired by [BO08, Corollary F.14]. It will turn out to be useful in the proof of Theorem D.
Lemma 2.3.
Let $\mathscr{R}$ be an equivalence relation on $(X,\nu)$. For every $i\in\{1,2\}$, let $(Z_{i},\nu_{i})$ be a standard probability space such that $(X,\nu)\cong(Z_{1}\times Z_{2},\nu_{1}\otimes\nu_{2})$. Let $\mathscr{S}$ be an ergodic equivalence relation on $(Z_{1},\nu_{1})$ such that $\mathscr{S}\times\operatorname{id}_{Z_{2}}\leq\mathscr{R}$ and $\mathscr{T}\leq\mathscr{R}$ a subequivalence relation.
If $\mathscr{S}\times\operatorname{id}_{Z_{2}}\npreceq_{\mathscr{R}}\mathscr{T}$, then there exists a hyperfinite ergodic subequivalence relation $\mathscr{S}_{0}\leq\mathscr{S}$ such that $\mathscr{S}_{0}\times\operatorname{id}_{Z_{2}}\npreceq_{\mathscr{R}}\mathscr{T}$.
Proof.
Choose a countable dense subset $\{g_{n}\mid n\geq 1\}$ in $[\mathscr{R}]$ with respect to the uniform metric $d_{u}$. Let $(f_{n})_{n\geq 1}$ be a dense sequence in $\operatorname{L}^{2}(Z_{1},\nu_{1})^{0}$. For any measurable subset $Z\subset Z_{1}$, we set $\nu_{Z}=\frac{1}{\nu_{1}(Z)}\nu_{1}|_{Z}$ and we let $\kappa^{Z}:[\mathscr{S}|_{Z}]\to\mathscr{U}(\operatorname{L}^{2}(Z,\nu_{Z})^{0})$ be the associated Koopman representation. The proof of the lemma will be concluded by constructing an increasing sequence $(\mathscr{S}_{n})_{n}$ of bounded subequivalence relations of $\mathscr{S}$ and a sequence $(\theta_{n})_{n}$ in $[\mathscr{S}]$ such that for all $n\geq 1$ and all $1\leq i,j\leq n$, we have $\theta_{n}\in[\mathscr{S}_{n}]$ and
$\varphi_{\mathscr{T}}(g_{i}(\theta_{n}\times\operatorname{id}_{Z_{2}})g_{j})\leq\frac{1}{n}\quad\text{and}\quad|\langle\kappa_{\theta_{n}}(f_{i}),f_{j}\rangle|\leq\frac{1}{n}.$
Let $\mathscr{S}_{1}\leq\mathscr{S}$ be the trivial subequivalence relation, $\theta_{1}=\operatorname{id}_{Z_{1}}$ and assume that $\mathscr{S}_{1},\dots,\mathscr{S}_{n-1}$ and $\theta_{1},\dots,\theta_{n-1}$ have been constructed. Since $\mathscr{S}_{n-1}$ is bounded, there exists a finite measurable partition $Z_{1}=X_{1}\sqcup\dots\sqcup X_{N}$ into $\mathscr{S}_{n-1}$-invariant sets such that $\mathscr{S}_{n-1}|_{X_{k}}$ is of type ${\rm I}_{t_{k}}$ for any $1\leq k\leq N$. For any $1\leq k\leq N$, let $Y_{k}\subset X_{k}$ be a fundamental domain for $\mathscr{S}_{n-1}|_{X_{k}}$ and let $\omega_{k}\in[\mathscr{S}_{n-1}|_{X_{k}}]$ such that $X_{k}=\sqcup_{r=0}^{t_{k}-1}\omega_{k}^{r}Y_{k}$.
Let $1\leq k\leq N$. Since $\mathscr{S}|_{Y_{k}}\times\operatorname{id}_{Z_{2}}\npreceq_{\mathscr{R}}\mathscr{T}$, Facts 2.2$(\rm iii)$ implies that there exists $\Psi_{k}\in[\mathscr{S}|_{Y_{k}}]$ such that for all $10\leq r\leq t_{k}-$ 1 and $1\leq i,j\leq n$, we have
(2.1) $\varphi_{\mathscr{T}}(g_{i}(\omega_{k}^{r}\Psi_{k}\omega_{k}^{-r}\times\operatorname{id}_{Z_{2}})g_{j})\leq\frac{1}{2nt_{k}N}\quad$
and
(2.2) $\quad|\langle\kappa^{Y_{k}}_{\Psi_{k}}((f_{i}\circ\omega_{k}^{r})|_{Y_{k}}),(f_{j}\circ\omega_{k}^{r})|_{Y_{k}}\rangle|\leq\frac{1}{n}.$
Next, by applying Facts 2.1$(\rm iii)$, there exist a measurable subset $\tilde{Y}_{k}\subset Y_{k}$ with $\nu_{1}(Y_{k}\setminus\tilde{Y}_{k})\leq\frac{1}{2nt_{k}N}$, a positive integer $p_{k}$ and $\tilde{\Psi}_{k}\in[\mathscr{S}|_{Y_{k}}]$ such that $\tilde{\Psi}_{k}=\Psi_{k}$ on $\tilde{Y}_{k}$ and $\tilde{\Psi}_{k}^{p_{k}}=\operatorname{id}_{Y_{k}}$. Fix $0\leq r\leq t_{k-1}$. Since $g_{i}(\omega_{k}^{r}\tilde{\Psi}_{k}\omega_{k}^{-r}\times\operatorname{id}_{Z_{2}})g_{j}=g_{i}(\omega_{k}^{r}\Psi_{k}\omega_{k}^{-r}\times\operatorname{id}_{Z_{2}})g_{j}$ on $g_{j}^{-1}(\omega_{k}^{r}\tilde{Y}_{k}\times Z_{2})$, it follows that for all $1\leq i,j\leq n$, we have
$\varphi_{\mathscr{T}}(g_{i}(\omega_{k}^{r}\tilde{\Psi}_{k}\omega_{k}^{-r}\times\operatorname{id}_{Z_{2}})g_{j})\leq\nu((Y_{k}\setminus\tilde{Y}_{k})\times Z_{2})+$
$\nu(\{x\in g_{j}^{-1}(\tilde{Y}_{k}\times Z_{2})\mid((g_{i}(\omega_{k}^{r}\Psi_{k}\omega_{k}^{-r}\times\operatorname{id}_{Z_{2}})g_{j})(x),x)\in\mathscr{T}\}),$
and thus, using (2.1) we have
(2.3) $\varphi_{\mathscr{T}}(g_{i}(\omega_{k}^{r}\tilde{\Psi}_{k}\omega_{k}^{-r}\times\operatorname{id}_{Z_{2}})g_{j})\leq\frac{1}{nt_{k}N}.$
Next, since $Y_{k}\subset X_{k}$ is a fundamental domain for $\mathscr{S}_{n-1}|_{X_{k}}$, we can extend $\tilde{\Psi}_{k}\in[\mathscr{S}|_{Y_{k}}]$ naturally to an element $\Omega_{k}\in[\mathscr{S}|_{X_{k}}]$. More precisely, we let $\Omega_{k}(\omega_{k}^{r}y)=\omega_{k}^{r}\tilde{\Psi}_{k}(y)$, for all $y\in Y_{k}$ and $10\leq r\leq t_{k}-$ 1. In this way, since $\nu_{X_{k}}(Y_{k})=\frac{1}{t_{k}}$, for all $1\leq i,j\leq n$, we deduce from (2.3) and (2.2) that
(2.4) $\varphi_{\mathscr{T}}(g_{i}(\Omega_{k}\times\operatorname{id}_{Z_{2}})g_{j})\leq\frac{1}{nN}\quad\text{and}\quad|\langle\kappa^{X_{k}}_{\Omega_{k}}(f_{i}|_{X_{k}}),f_{j}|_{X_{k}}\rangle|\leq\frac{1}{n}.$
Indeed, note that (2.3) implies that
$\varphi_{\mathscr{T}}(g_{i}(\Omega_{k}\times\operatorname{id}_{Z_{2}})g_{j})$
$=\sum_{r=0}^{t_{k}-1}\nu(\{x\in g_{j}^{-1}(\omega_{k}^{r}Y_{k}\times Z_{2})\mid((g_{i}(\Omega_{k}\times\operatorname{id}_{Z_{2}})g_{j})(x),x)\in\mathscr{T}\})$
$=\sum_{r=0}^{t_{k}-1}\nu(\{x\in g_{j}^{-1}(\omega_{k}^{r}Y_{k}\times Z_{2})\mid((g_{i}(\omega_{k}^{r}\tilde{\Psi}_{k}\omega_{k}^{-r}\times\operatorname{id}_{Z_{2}})g_{j})(x),x)\in\mathscr{T}\})$
$\leq\sum_{r=0}^{t_{k}-1}\varphi_{\mathscr{T}}(g_{i}(\omega_{k}^{r}\tilde{\Psi}_{k}\omega_{k}^{-r}\times\operatorname{id}_{Z_{2}})g_{j})\leq\frac{1}{nN}.$
The second part of (2.4) follows in a similar way and we omit the details.
Next, define $\theta_{n}\in\operatorname{Aut}(Z_{1},\nu_{1})$ by $\theta_{n}=\Omega_{k}$ on $X_{k}$ for any $1\leq k\leq N$. By construction, we deduce that $\theta_{n}\in[\mathscr{S}]$ and that there exists an integer $p$ (e.g. take $p=p_{1}\cdots p_{N}$) for which $\theta_{n}^{p}=\operatorname{id}_{Z_{1}}$. Note that for all $1\leq i,j\leq n$, it follows from (2.4) that
$\varphi_{\mathscr{T}}(g_{i}(\theta_{n}\times\operatorname{id}_{Z_{2}})g_{j})$
$=\nu(\bigcup_{k=1}^{N}\{x\in g_{j}^{-1}(X_{k}\times Z_{2})\mid((g_{i}(\theta_{n}\times\operatorname{id}_{Z_{2}})g_{j})(x),x)\in\mathscr{T}\})$
$\leq\sum_{k=1}^{N}\nu(\{x\in g_{j}^{-1}(X_{k}\times Z_{2})\mid((g_{i}(\Omega_{k}\times\operatorname{id}_{Z_{2}})g_{j})(x),x)\in\mathscr{T}\})$
$=\sum_{k=1}^{N}\varphi_{\mathscr{T}}(g_{i}(\Omega_{k}\times\operatorname{id}_{Z_{2}})g_{j})\leq\sum_{k=1}^{N}\frac{1}{nN}=\frac{1}{n}.$
We derive in a similar way using (2.4) that for all $1\leq i,j\leq n$, we have
(2.5) $|\langle\kappa^{Z_{1}}_{\theta_{n}}(f_{i}),f_{j}\rangle|\leq\frac{1}{n}.$
We can now define $\mathscr{S}_{n}$ as the subequivalence relation of $\mathscr{S}$ generated by $\mathscr{S}_{n-1}$ and $\theta_{n}$. Since $\theta_{n}|_{X_{k}}$ commutes with $\omega_{k}$ for any $1\leq k\leq N$, one can deduce that $\mathscr{S}_{n}$ is bounded. By letting $\mathscr{S}_{0}=\bigcup_{k\geq 1}\mathscr{S}_{k}$, we derive that $\mathscr{S}_{0}$ is hyperfinite and that $\mathscr{S}_{0}\times\operatorname{id}_{Z_{2}}\npreceq_{\mathscr{R}}\mathscr{T}$. By (2.5), the Koopman representation associated with $[\mathscr{S}_{0}]$ is weakly mixing. This implies by Facts 2.1(i) that $\mathscr{S}_{0}$ is ergodic. ∎
The notion of strong primeness for type ${\rm II_{1}}$ factors was introduced by Isono [Is16]. We adapt this notion to the setting of ergodic equivalence relations.
Definition 2.4.
Let $\mathscr{R}$ be an ergodic equivalence relation. We say that $\mathscr{R}$ is strongly prime if for any ergodic equivalence relations $\mathscr{S}$, $\mathscr{T}_{1}$, $\mathscr{T}_{2}$ for which ×$\mathscr{R}\times\mathscr{S}\cong\mathscr{T}_{1}\times\mathscr{T}_{2}$, then there exists $i\in\{1,2\}$ such that we have $\mathscr{T}_{i}\preceq_{\mathscr{R}\times\mathscr{S}}\mathscr{S}$ as subequivalence relations.
2.3. Algebraic groups
Let $k$ be a local field, that is, a nondiscrete locally compact field. Throughout, we assume that the characteristic of $k$ is zero. Then $k$ is $\mathbb{R}$, $\mathbb{C}$ or a finite extension of $\mathbb{Q}_{p}$ for some prime $p$. In this paper, by a $k$-group $\mathbf{G}$, we simply mean a linear algebraic group defined over $k$. We denote by $\mathbf{G}^{0}$ the Zariski connected component of the identity element $e\in\mathbf{G}$. We say that $\mathbf{G}$ is Zariski connected if $\mathbf{G}=\mathbf{G}^{0}$. We say that $\mathbf{G}$ is almost $k$-simple if $\mathbf{G}$ is not abelian and the only proper normal $k$-closed subgroups are finite. In that case, since $\mathbf{G}$ is connected, finite normal subgroups are contained in $\mathscr{Z}(\mathbf{G})$. We say that $\mathbf{G}$ is $k$-simple if $\mathbf{G}$ is not abelian and the only normal $k$-closed subgroups are $\{e\}$ and $\mathbf{G}$. If $\mathbf{G}$ is (almost) $k$-simple, then we moreover say that $\mathbf{G}$ is $k$-isotropic if $\operatorname{rk}_{k}(\mathbf{G})\geq 1$. We denote by $G=\mathbf{G}(k)$ the locally compact second countable group of its $k$-points. Then $G$ is a $k$-analytic Lie group and we naturally have $\operatorname{Lie}(\mathbf{G})(k)=\operatorname{Lie}(G)$ as $k$-Lie algebras.
By a $k$-$\mathbf{G}$-variety $\mathbf{V}$, we simply mean an algebraic variety defined over $k$ that is endowed with an algebraic action $\mathbf{G}\curvearrowright\mathbf{V}$ so that the action map $\mathbf{G}\times\mathbf{V}\to\mathbf{V}:(g,v)\mapsto gv$ is a $k$-morphism. We denote by $V=\mathbf{V}(k)$ the Hausdorff locally compact second countable topological space of its $k$-points. If $\mathbf{V}$ is smooth, then $V$ is a $k$-analytic manifold. In the case $\mathbf{V}=\mathbf{G}/\mathbf{H}$, where $\mathbf{H}<\mathbf{G}$ is a $k$-subgroup, we denote by $\pi:\mathbf{G}\to\mathbf{G}/\mathbf{H}$ the $\mathbf{G}$-equivariant $k$-morphism such that $\pi(e)=\mathbf{H}$. Then $v=\pi(e)\in(\mathbf{G}/\mathbf{H})(k)$ and $H=\mathbf{H}(k)=\operatorname{Stab}_{G}(v)$.
Facts 2.5.
Keep the same notation as above. We will be using the following useful properties.
  • $(\rm i)$
    The continuous action $G\curvearrowright V$ has locally closed orbits for the $k$-analytic topology and so the quotient space
    $G\backslash V=\left\{Gv\mid v\in V\right\}$
    is a standard Borel space.
  • $(\rm ii)$
    The continuous action $G\curvearrowright\operatorname{Prob}(V)$ has locally closed orbits and so the quotient space
    $G\backslash\operatorname{Prob}(V)=\left\{G\mu\mid\mu\in\operatorname{Prob}(V)\right\}$
    is a standard Borel space.
  • $(\rm iii)$
    Assume that $\mathbf{V}=\mathbf{G}/\mathbf{H}$, where $\mathbf{H}<\mathbf{G}$ is a $k$-subgroup. Set $v=\pi(e)\in(\mathbf{G}/\mathbf{H})(k)$ and $H=\mathbf{H}(k)=\operatorname{Stab}_{G}(v)$. The orbit map $G/H\to Gv:gH\to gv$ is a homeomorphism, we may identify the $G$-orbit $Gv$ with $G/H$ and we may regard $G/H\subset(\mathbf{G}/\mathbf{H})(k)$ as a closed and open subset.
Proof.
$(\rm i)$ This follows from [Zi84, Theorem 2.1.14, Proposition 3.1.3].
$(\rm ii)$ This follows from [Zi84, Theorem 2.1.14, Proposition 3.2.6] and more generally [BDL14, Theorem 1.7].
$(\rm iii)$ This follows from [Ma91, Proposition I.2.1.4]. ∎
2.4. Algebraic representations of equivalence relations
Let $\mathscr{R}$ be an equivalence relation on a standard probability space $(X,\nu)$, $H$ a locally compact second countable group and $\alpha:\mathscr{R}\to H$ a measurable $1$-cocycle. Let $Z$ be a standard Borel space endowed with a Borel action $H\curvearrowright Z$ and $f:X\to Z$ a Borel map. We say that $f:X\to Z$ is $(\mathscr{R},\alpha)$-equivariant if for $m$-almost every $(x,y)\in\mathscr{R}$, we have $f(x)=\alpha(x,y)f(y)$. We say that $f:X\to Z$ is $\mathscr{R}$-invariant if for $m$-almost every $(x,y)\in\mathscr{R}$, we have $f(x)=f(y)$.
The following dichotomy theorem generalizes Zimmer’s result [Zi84, Theorem 9.2.3] (see also [BDL14, Theorem 6.1] for the case of measurable $1$-cocycles associated with nonsingular group actions). For the sake of completeness, we give a short proof.
Theorem 2.6.
Let $k$ be a local field of characteristic zero and $\mathbf{G}$ a $k$-isotropic almost $k$-simple algebraic $k$-group. Let $\mathscr{R}$ be an amenable ergodic equivalence relation on $(X,\nu)$ and $\alpha:\mathscr{R}\to G$ a measurable $1$-cocycle. Then at least one of the following assertions holds:
  • $(\rm i)$
    There exist a proper $k$-subgroup $\mathbf{H}<\mathbf{G}$ and an $(\mathscr{R},\alpha)$-equivariant measurable map $\psi:X\to G/H$, where we regard $G/H\subset(\mathbf{G}/\mathbf{H})(k)$ with $H=\mathbf{H}(k)$.
  • $(\rm ii)$
    There exist a compact subgroup $L<G$ and an $(\mathscr{R},\alpha)$-equivariant measurable map $f:X\to G/L$.
Observe that in the case $k=\mathbb{R}$, assertion $(\rm i)$ always holds.
Proof.
Choose a minimal parabolic $k$-subgroup $\mathbf{P}<\mathbf{G}$ and denote by $\mathbf{V}=\mathbf{G}/\mathbf{P}$ the homogeneous $k$-$\mathbf{G}$-variety. Write $G=\mathbf{G}(k)$, $P=\mathbf{P}(k)$ and $V=(\mathbf{G}/\mathbf{P})(k)=\mathbf{G}(k)/\mathbf{P}(k)$ (see [Bo91, Proposition 20.5]). Since $V$ is compact and since $\mathscr{R}$ is amenable, there exists an $(\mathscr{R},\alpha)$-equivariant measurable map $\beta:X\to\operatorname{Prob}(V)$ (see [Zi84, Proposition 4.3.9]). By Facts 2.5$(\rm ii)$, the continuous action $G\curvearrowright\operatorname{Prob}(V)$ has locally closed orbits and so the quotient space $G\backslash\operatorname{Prob}(V)$ is a standard Borel space. Denote by $p:\operatorname{Prob}(V)\to G\backslash\operatorname{Prob}(V)$ the quotient Borel map. Then the measurable map $p\circ\beta:X\to G\backslash\operatorname{Prob}(V)$ is $\mathscr{R}$-invariant. Since $\mathscr{R}$ is ergodic, $p\circ\beta:X\to G\backslash\operatorname{Prob}(V)$ is $\nu$-almost everywhere constant and so there exists $\mu\in\operatorname{Prob}(V)$ such that $\beta(X)$ is essentially contained in $G\mu$. Set $L=\operatorname{Stab}_{G}(\mu)<G$. By [Zi84, Theorem 2.1.14], the orbit map $G/L\to G\mu:gL\to g\mu$ is a homeomorphism and so we may regard $\beta:X\to G/L$ as an $(\mathscr{R},\alpha)$-equivariant measurable map.
Denote by $\mathbf{H}$ the Zariski closure of $L$ in $\mathbf{G}$. Then $\mathbf{H}<\mathbf{G}$ is a $k$-subgroup and we set $H=\mathbf{H}(k)$. By [BDL14, Proposition 1.9], there exists a $k$-subgroup $\mathbf{H}_{0}<\mathbf{H}<\mathbf{G}$ such that $\mathbf{H}_{0}\lhd\mathbf{H}$ is normal, the image of $L$ is precompact in $(\mathbf{H}/\mathbf{H}_{0})(k)$ and $\mu$ is supported on $\mathbf{V}^{\mathbf{H}_{0}}\cap V$. Since $\mathbf{V}^{\mathbf{G}}=\emptyset$, we have $\mathbf{H}_{0}\neq\mathbf{G}$.
Assume that $\mathbf{H}_{0}$ is infinite. Since $\mathbf{H}_{0}\lhd\mathbf{H}$, $\mathbf{H}_{0}\neq\mathbf{G}$ and $\mathbf{G}$ is almost $k$-simple, we have $\mathbf{H}\neq\mathbf{G}$. Since $L<H$, we may consider the $G$-equivariant measurable factor map $q:G/L\to G/H$. Then the measurable map $f=q\circ\beta:X\to G/H$ is $(\mathscr{R},\alpha)$-equivariant.
Assume that $\mathbf{H}_{0}$ is finite. Set $H_{0}=\mathbf{H}_{0}(k)$. The image of $L$ in $(\mathbf{H}/\mathbf{H}_{0})(k)$ is contained in $H/H_{0}\subset(\mathbf{H}/\mathbf{H}_{0})(k)$, which is closed in $(\mathbf{H}/\mathbf{H}_{0})(k)$. Then the image of $L$ in $H/H_{0}$ is precompact. Since $H_{0}<H$ is finite, this further implies that $L<H$ is compact. ∎
We use [BDL14, Theorem 6.1] in combination with [TD14, Theorem 14] to show that Zariski dense discrete subgroups are not inner amenable.
Proposition 2.7.
Let $k$ be a local field of characteristic zero and $\mathbf{G}$ a $k$-isotropic almost $k$-simple algebraic $k$-group. Let $\Gamma<\mathbf{G}(k)$ be a Zariski dense discrete subgroup. Then $\Gamma$ is not inner amenable.
Proof.
By contradiction, assume that $\Gamma<\mathbf{G}(k)$ is inner amenable. By [TD14, Theorem 14], there exists a short exact sequence $1\to N\to\Gamma\to K\to 1$, where $K$ is amenable and either
  • $(\rm i)$
    $\mathscr{Z}(N)$ is infinite, or
  • $(\rm ii)$
    $N=LM$, where $L$ and $M$ are commuting normal subgroups of $\Gamma$ such that $M$ is infinite and amenable, and $L\cap M$ is finite.
In case $(\rm i)$, since $\mathscr{Z}(N)\lhd\Gamma$ is an infinite normal subgroup, it follows that $\mathscr{Z}(N)$ is Zariski dense in $\mathbf{G}$. This would imply that $\mathbf{G}$ is abelian, which is absurd.
In case $(\rm ii)$, since $M\lhd\Gamma$ is an infinite normal subgroup, it follows that $M$ is Zariski dense in $\mathbf{G}$. Since $M$ is amenable, applying [BDL14, Theorem 6.1] to the trivial action $M\curvearrowright\{\bullet\}$ which is amenable, at least one of the following assertions holds:
  • There exist a proper $k$-subgroup $\mathbf{H}<\mathbf{G}$ and an $M$-equivariant measurable map $\psi:\{\bullet\}\to G/H$, where we regard $G/H\subset(\mathbf{G}/\mathbf{H})(k)$ with $H=\mathbf{H}(k)$. Then there exists $g\in G$ such that $M<gHg^{-1}$ and so the Zariski closure of $M$ is a proper $k$-subgroup of $\mathbf{G}$, which is a contradiction.
  • There exist a compact subgroup $L<G$ and an $M$-equivariant measurable map $f:\{\bullet\}\to G/L$. Then there exists $g\in G$ such that $M<gLg^{-1}$ and so $M$ is compact, which is a contradiction.
This shows that $\Gamma$ is not inner amenable. ∎
Remark 2.8.
Let $k$ be a local field of characteristic zero and $\mathbf{G}$ a $k$-isotropic $k$-simple algebraic $k$-group. Let $\Gamma<\mathbf{G}(k)$ be a Zariski dense discrete subgroup. Then $\Gamma$ is icc meaning that $\Gamma$ has infinite conjugacy classes. Indeed, let $\gamma\in\Gamma$ be such that its conjugacy class $C(\gamma)=\left\{h\gamma h^{-1}\mid h\in\Gamma\right\}$ is finite. Then $\mathscr{Z}_{\Gamma}(\gamma)<\Gamma$ has finite index and so $\mathscr{Z}_{\Gamma}(\gamma)$ is Zariski dense in $\mathbf{G}$ by Borel’s density theorem [Bo60]. This implies that $\gamma\in\mathscr{Z}(\mathbf{G})$. Since $\mathbf{G}$ is $k$-simple, we have $\gamma=e$.
2.5. Chabauty topology and Wijsman topology
Let $G$ be a locally compact second countable group. Consider the space $\operatorname{Sub}(G)$ of all closed subgroups of $G$. Endowed with the Chabauty topology, $\operatorname{Sub}(G)$ is a compact metrizable space and the conjugation action $G\curvearrowright\operatorname{Sub}(G)$ is continuous (see [Ch50]). Recall that for any sequence $(H_{n})_{n\in\mathbb{N}}$ in $\operatorname{Sub}(G)$ and any element $H\in\operatorname{Sub}(G)$, we have $H_{n}\to H$ in $\operatorname{Sub}(G)$ if and only if the following two conditions hold:
  • $(\rm i)$
    For every $x\in H$, there exists a sequence $(x_{n})_{n\in\mathbb{N}}$ in $G$ such that $x_{n}\in H_{n}$ for every $n\in\mathbb{N}$ and $x_{n}\to x$ in $G$.
  • $(\rm ii)$
    For every increasing sequence $(n_{j})_{j\in\mathbb{N}}$ in $\mathbb{N}$, every sequence $(x_{j})_{j\in\mathbb{N}}$ in $G$ such that $x_{j}\in H_{n_{j}}$ for every $j\in\mathbb{N}$, and every element $x\in G$, if $x_{j}\to x$ in $G$, then $x\in H$.
We refer the reader to [BP92, Section E.1] for further details. Then for every $K\in\operatorname{Sub}(G)$, the subset
(2.6) $\left\{H\in\operatorname{Sub}(G)\mid H<K\right\}\subset\operatorname{Sub}(G)\quad\text{is closed}.$
Let $\mathscr{G}$ be a Polish group and choose a compatible complete metric $d:\mathscr{G}\times\mathscr{G}\to\mathbb{R}_{+}$. Denote by $\operatorname{CL}(\mathscr{G})$ the space of all nonempty closed subsets of $\mathscr{G}$. For all $x\in\mathscr{G}$, $A\in\operatorname{CL}(\mathscr{G})$, set $d(x,A)=\inf\left\{d(x,a)\mid a\in A\right\}$. Following [Be90], the Wijsman topology $\tau_{W(d)}$ on $\operatorname{CL}(\mathscr{G})$ is defined as the weakest topology on $\operatorname{CL}(\mathscr{G})$ that makes the maps $\operatorname{CL}(\mathscr{G})\to\mathbb{R}_{+}:A\mapsto d(x,A)$ continuous for all $x\in\mathscr{G}$. As explained in [Be90], the Borel structure induced by the Wijsman topology $\tau_{W(d)}$ coincides with the Effros-Borel structure on $\operatorname{CL}(\mathscr{G})$. By [Be90, Theorem 4.3], $(\operatorname{CL}(\mathscr{G}),\tau_{W(d)})$ is a Polish space. Consider the space $\operatorname{Sub}(\mathscr{G})$ of all closed subgroups of $\mathscr{G}$. Then $\operatorname{Sub}(\mathscr{G})\subset\operatorname{CL}(\mathscr{G})$ is closed with respect to $\tau_{W(d)}$ and so $(\operatorname{Sub}(\mathscr{G}),\tau_{W(d)})$ is a Polish space. Observe that the subset
(2.7) $\left\{(h,\mathscr{H})\in\mathscr{G}\times\operatorname{Sub}(\mathscr{G})\mid h\in\mathscr{H}\right\}\subset\mathscr{G}\times\operatorname{Sub}(\mathscr{G})\quad\text{is closed}.$
When $\mathscr{G}=G$ is a locally compact second countable group and $d_{G}:G\times G\to\mathbb{R}_{+}$ is a left invariant compatible proper complete metric (see [St73]), the Wijsman topology $\tau_{W(d_{G})}$ and the Chabauty topology coincide on $\operatorname{Sub}(G)$ (see [BLLN89, Section 5]).
Let now $G$ be a locally compact second countable group. Choose a left invariant compatible proper complete metric $d_{G}:G\times G\to\mathbb{R}_{+}$ (see [St73]). Then $d=\min(d_{G},1):G\times G\to\mathbb{R}_{+}$ is a left invariant compatible complete metric such that $d\leq 1$. Let $(X,\nu)$ be a standard probability space. Denote by $\mathscr{G}=\operatorname{L}^{0}(X,\nu,G)$ the space of all $\nu$-equivalence classes of measurable maps $f:X\to G$. Endowed with the topology of convergence in measure, $\mathscr{G}$ is a Polish group (see e.g. [Ke10, Section 19]). More precisely, define the metric $d_{\mathscr{G}}:\mathscr{G}\times\mathscr{G}\to\mathbb{R}_{+}$ by the formula
$\forall f_{1},f_{2}\in\mathscr{G},\quad d_{\mathscr{G}}(f_{1},f_{2})=\int_{X}d(f_{1}(x),f_{2}(x))\,{\rm d}\nu(x).$
Then $d_{\mathscr{G}}$ is a left invariant compatible complete metric on $\mathscr{G}$. For every $H\in\operatorname{Sub}(G)$, we have that $\mathscr{H}=\operatorname{L}^{0}(X,\nu,H)\in\operatorname{Sub}(\mathscr{G})$. The following technical result will be very useful in the proof of Theorem 3.3.
Proposition 2.9.
Keep the same notation as above. Then the map
$\Psi:\operatorname{Sub}(G)\to\operatorname{Sub}(\mathscr{G}):H\mapsto\operatorname{L}^{0}(X,\nu,H)$
is Borel.
Proof.
Consider the Wijsman topology $\tau=\tau_{W(d_{\mathscr{G}})}$ on $\operatorname{Sub}(\mathscr{G})$. Fix a countable dense subset $D\subset\mathscr{G}$. For every $f\in D$ and all $\alpha,\beta\in\mathbb{Q}\cap[0,1]$ such that $\alpha<\beta$, define the open subsets
$O(f,\alpha)^{+}$ =$=\left\{\mathscr{K}\in\operatorname{Sub}(\mathscr{G})\mid d_{\mathscr{G}}(f,\mathscr{K})>\alpha\right\}$
$O(f,\beta)^{-}$ =$=\left\{\mathscr{K}\in\operatorname{Sub}(\mathscr{G})\mid d_{\mathscr{G}}(f,\mathscr{K})<\beta\right\}$
$O(f,\alpha,\beta)$ $=O(f,\alpha)^{+}\cap O(f,\beta)^{-}.$
Then the countable set $\left\{O(f,\alpha,\beta)\mid f\in D,\alpha,\beta\in\mathbb{Q}\cap[0,1],\alpha<\beta\right\}$ is a subbase for the topology $\tau$ on $\operatorname{Sub}(\mathscr{G})$. Thus, in order to show that the map $\Psi:\operatorname{Sub}(G)\to\operatorname{Sub}(\mathscr{G})$ is Borel, it suffices to show that $\Psi^{-1}(O(f,\alpha)^{+})$ and $\Psi^{-1}(O(f,\alpha)^{-})$ are Borel subsets of $\operatorname{Sub}(G)$ for every $f\in D$ and every $\alpha\in\mathbb{Q}\cap[0,1]$.
Claim 2.10.
Let $(H_{n})_{n\in\mathbb{N}}$ be a sequence in $\operatorname{Sub}(G)$ and $H\in\operatorname{Sub}(G)$ such that $H_{n}\to H$ in $\operatorname{Sub}(G)$. Set $\mathscr{H}_{n}=\operatorname{L}^{0}(X,\nu,H_{n})$ for every $n\in\mathbb{N}$ and $\mathscr{H}=\operatorname{L}^{0}(X,\nu,H)$. Then for every $f\in\mathscr{H}$, there exists a sequence $(f_{n})_{n\in\mathbb{N}}$ in $\mathscr{G}$ such that $f_{n}\in\mathscr{H}_{n}$ for every $n\in\mathbb{N}$ and $\lim_{n}d_{\mathscr{G}}(f,f_{n})=0$.
Indeed, let $f\in\mathscr{H}$. Without loss of generality, we may assume that $X$ is a compact metrizable space and $\nu\in\operatorname{Prob}(X)$ is a Borel probability measure. Regard $f:X\to H$ as a measurable function. By Lusin’s theorem, for every $n\geq 1$, there exists a closed subset $X_{n}\subset X$ such that $\nu(X_{n})\geq 1-1/n$ and $f|_{X_{n}}:X_{n}\to H$ is continuous. By compactness, there exist $r\geq 1$ and $h_{1},\dots,h_{r}\in H$ such that $X_{n}\subset\bigcup_{j=1}^{r}f^{-1}(B_{d}(h_{j},1/n))$, where $B_{d}(h,\varepsilon)\subset G$ denotes the open ball in $G$ of center $h\in G$ and radius $\varepsilon>0$. Set $Y_{1}=X_{n}\cap f^{-1}(B_{d}(h_{1},1/n))$ and $Y_{j}=X_{n}\cap f^{-1}(B_{d}(h_{j},1/n))\setminus\bigcup_{i=1}^{j-1}Y_{i}$ for every $2\leq j\leq r$. Then $(Y_{j})_{1\leq j\leq r}$ is a measurable partition of $X_{n}$. Since $H_{n}\to H$ in $\operatorname{Sub}(G)$, for every $1\leq j\leq r$, we may find $h_{j}^{n}\in B_{d}(h_{j},1/n)\cap H_{n}$. Define the measurable map $f_{n}:X\to H_{n}$ by $f_{n}|_{X\setminus X_{n}}=e\mathbf{1}_{X\setminus X_{n}}$ and $f_{n}|_{Y_{j}}=h_{j}^{n}\mathbf{1}_{Y_{j}}$ for every $1\leq j\leq r$. Then we have
$\limsup_{n}d_{\mathscr{G}}(f,f_{n})$ $=\limsup_{n}\int_{X}d(f(x),f_{n}(x))\,{\rm d}\nu(x)$
$\leq\limsup_{n}\left(\nu(X\setminus X_{n})+\frac{2}{n}\nu(X_{n})\right)=0.$
This finishes the proof of Claim 2.10.
Claim 2.11.
Let $(H_{n})_{n\in\mathbb{N}}$ be a sequence in $\operatorname{Sub}(G)$ and $H\in\operatorname{Sub}(G)$ such that $H_{n}\to H$ in $\operatorname{Sub}(G)$. Set $\mathscr{H}_{n}=\operatorname{L}^{0}(X,\nu,H_{n})$ for every $n\in\mathbb{N}$ and $\mathscr{H}=\operatorname{L}^{0}(X,\nu,H)$. Then for every $f\in\mathscr{G}$, we have
$\limsup_{n}d_{\mathscr{G}}(f,\mathscr{H}_{n})\leq d_{\mathscr{G}}(f,\mathscr{H}).$
Indeed, set $\alpha=d_{\mathscr{G}}(f,\mathscr{H})$ and $\beta=\limsup_{n}d_{\mathscr{G}}(f,\mathscr{H}_{n})$. Choose an increasing sequence $(n_{j})_{j\in\mathbb{N}}$ in $\mathbb{N}^{*}$ such that $\beta=\lim_{j}d_{\mathscr{G}}(f,\mathscr{H}_{n_{j}})$. Choose a sequence $(f_{n})_{n\in\mathbb{N}}$ in $\mathscr{H}$ such that $\alpha\leq d_{\mathscr{G}}(f,f_{n})<\alpha+1/n$ for every $n\geq 1$. By Claim 2.10, we may choose a sequence $(h_{n})_{n\geq 1}$ in $\mathscr{G}$ such that $h_{n}\in\mathscr{H}_{n}$ and $d_{\mathscr{G}}(f_{n},h_{n})\leq 1/n$ for every $n\geq 1$. In particular, for every $j\in\mathbb{N}$, we have
$d_{\mathscr{G}}(f,\mathscr{H}_{n_{j}})\leq d_{\mathscr{G}}(f,h_{n_{j}})\leq d_{\mathscr{G}}(f,f_{n_{j}})+d_{\mathscr{G}}(f_{n_{j}},h_{n_{j}})\leq\alpha+\frac{2}{n_{j}}.$
This implies that $\beta\leq\alpha$ and finishes the proof of Claim 2.11.
Let $f\in D$ and $\alpha\in\mathbb{Q}\cap[0,1]$. Claim 2.11 implies that $\Psi^{-1}(O(f,\alpha)^{-})\subset\operatorname{Sub}(G)$ is open. Observe that
$O(f,\alpha)^{+}=\bigcup_{j=1}^{\infty}\left\{\mathscr{K}\in\operatorname{Sub}(\mathscr{G})\mid d_{\mathscr{G}}(f,\mathscr{K})\geq\alpha+1/j\right\}.$
Claim 2.11 implies that $\Psi^{-1}(O(f,\alpha)^{+})\subset\operatorname{Sub}(G)$ is a countable union of closed subsets and so $\Psi^{-1}(O(f,\alpha)^{+})\subset\operatorname{Sub}(G)$ is Borel. Therefore, we have showed that the map $\Psi:\operatorname{Sub}(G)\to\operatorname{Sub}(\mathscr{G})$ is Borel. ∎
3. Proofs of the main results
We will frequently use the following version of the Jankov–von Neumann selection theorem.
Theorem 3.1.
Let $(X,\nu)$ be a standard probability space, $Y$ a Polish space and $B\subset X\times Y$ a Borel subset such that $\pi_{X}(B)=X$, where $\pi_{X}:X\times Y\to X$ is the projection map. Then there exists a Borel map $f:X\to Y$ such that $(x,f(x))\in B$ for $\nu$-almost every $x\in X$.
Proof.
By the Jankov–von Neumann selection theorem, there exists a $\sigma(\Sigma_{1}^{1})$-measurable map $f:X\to Y$ such that $(x,f(x))$ for every $x\in X$ (see [Ke95, Theorem 18.1]). We then use two well-known facts. Firstly, any $\sigma(\Sigma_{1}^{1})$-measurable set is Lebesgue measurable (see [Ke95, Theorem 29.7]). Secondly, any Lebesgue measurable map $X\to Y$ coincides $\nu$-almost everywhere with a Borel map. Therefore, upon modifying $f:X\to Y$ on a $\nu$-null Borel subset, we may assume that the map $f:X\to Y$ is Borel and satisfies $(x,f(x))\in B$ for $\nu$-almost every $x\in X$. ∎
Next, let $(X,\nu)$ and $(Y,\eta)$ be two standard probability spaces. Recall that the spaces $\operatorname{L}^{0}(X\times Y,\nu\otimes\eta,\mathbb{C})$, $\operatorname{L}^{0}(Y,\eta,\mathbb{C})$ and $\operatorname{L}^{0}(X,\nu,\operatorname{L}^{0}(Y,\eta,\mathbb{C}))$ are Polish spaces when endowed with the topology of convergence in measure. Moreover, the well-defined map
$\iota:\operatorname{L}^{0}(X\times Y,\nu\otimes\eta,\mathbb{C})\to\operatorname{L}^{0}(X,\nu,\operatorname{L}^{0}(Y,\eta,\mathbb{C})):f\mapsto(x\mapsto f(x,\cdot\,))$
is isometric (hence injective) and surjective. Therefore, we may and will identify $\operatorname{L}^{0}(X\times Y,\nu\otimes\eta,\mathbb{C})$ and $\operatorname{L}^{0}(X,\nu,\operatorname{L}^{0}(Y,\eta,\mathbb{C}))$.
3.1. Zimmer’s method for product equivalence relations
Let $k$ be a local field of characteristic zero and $\mathbf{G}$ a $k$-group. Set $G=\mathbf{G}(k)$.
Let $\mathscr{U}$ be an ergodic equivalence relation on $(U,\nu)$ and $\mathscr{V}$ an ergodic equivalence relation on $(V,\eta)$. Set $\mathscr{W}=\mathscr{U}\times\mathscr{V}$ and $(W,\zeta)=(U\times V,\nu\otimes\eta)$. Let $\alpha:\mathscr{W}\to G$ be a Borel $1$-cocycle. Following [Zi81], we define the set $\mathfrak{A}(\alpha)$ that consists of all pairs $(\varphi,\mathfrak{H})$, where $\varphi:W\to G$ is a Borel map, $\mathfrak{H}:U\to\operatorname{Sub}(G):u\mapsto H_{u}$ is a Borel map such that for $\nu$-almost every $u\in U$, there exists a $k$-subgroup $\mathbf{H}_{u}<\mathbf{G}$ such that $\mathbf{H}_{u}(k)=H_{u}$ and the Borel map $V\to G/H_{u}:v\mapsto\varphi(u,v)H_{u}$ is $(\mathscr{V},\alpha_{u})$-equivariant, where $\alpha_{u}:\mathscr{V}\to G:(v_{1},v_{2})\mapsto\alpha((u,v_{1}),(u,v_{2}))$. Observe that the trivial pair $(\varphi,\mathfrak{H})$ defined by $\varphi:W\to G:(u,v)\mapsto e$ and $\mathfrak{H}:U\to\operatorname{Sub}(G):u\mapsto G$ belongs to $\mathfrak{A}(\alpha)$. Thus, $\mathfrak{A}(\alpha)$ is not the empty set.
As explained in [Zi81], the set $\mathfrak{A}(\alpha)$ is stable under the following cutting and pasting procedures. Firstly, let $(\varphi,\mathfrak{H})\in\mathfrak{A}(\alpha)$ and $\lambda\in[\mathscr{U}]$. Define the pair $(\psi,\mathfrak{J})=\lambda_{\ast}(\varphi,\mathfrak{H})$ by $\psi:W\to G:(u,v)\mapsto\alpha((u,v),(\lambda u,v))\varphi(\lambda u,v)$ and $\mathfrak{J}:U\to\operatorname{Sub}(G):u\mapsto H_{\lambda u}$. Then a straightforward computation shows that $(\psi,\mathfrak{J})\in\mathfrak{A}(\alpha)$. Secondly, let $(\varphi^{i},\mathfrak{H}^{i})\in\mathfrak{A}(\alpha)$ for every $i\in\{1,2\}$. Let $U=U_{1}\sqcup U_{2}$ be a measurable partition. Define the pair $(\varphi,\mathfrak{H})$ by $\varphi|_{U_{i}\times V}=\varphi^{i}|_{U_{i}\times V}$ and $\mathfrak{H}|_{U_{i}\times V}=\mathfrak{H}^{i}|_{U_{i}\times V}$ for every $i\in\{1,2\}$. Then it is plain to see that $(\varphi,\mathfrak{H})\in\mathfrak{A}(\alpha)$. We record the following useful result.
Lemma 3.2.
There exists a pair $(\varphi,\mathfrak{H})\in\mathfrak{A}(\alpha)$ such that the $\nu$-almost everywhere defined map $U\mapsto\mathbb{N}:u\mapsto\dim(\mathbf{H}_{u})$ is $\nu$-almost everywhere constant and for any other pair $(\psi,\mathfrak{J})\in\mathfrak{A}(\alpha)$, we have $\dim(\mathbf{J}_{u})\geq\dim(\mathbf{H}_{u})$ for $\nu$-almost every $u\in U$.
Proof.
Consider the nonempty finite subset $\mathscr{F}\subset\mathbb{N}$ that consists of all integers $j\in\mathbb{N}$ for which there exists $(\psi,\mathfrak{J})\in\mathfrak{A}(\alpha)$ such that $\nu(\{u\in U\mid\dim(\mathbf{J}_{u})=j\})>0$. Set $d=\min(\mathscr{F})$ and choose a pair $(\psi,\mathfrak{J})\in\mathfrak{A}(\alpha)$ such that $\nu(\{u\in U\mid\dim(\mathbf{J}_{u})=d\})>0$. Choose $n\geq 1$ and a measurable subset $U_{1}\subset\{u\in U\mid\dim(\mathbf{J}_{u})=d\}$ such that $\nu(U_{1})=\frac{1}{n}$. Choose measurable subsets $U_{2},\dots,U_{n}\subset X$ so that $(U_{i})_{1\leq i\leq n}$ forms a measurable partition of $X$. Since $\mathscr{U}$ is ergodic on $(U,\nu)$, for every $1\leq i\leq n$, we may choose $\lambda_{i}\in[\mathscr{U}]$ such that $\lambda_{i}(U_{i})=U_{1}$. Then for every $1\leq i\leq n$, we have $(\psi^{i},\mathfrak{J}^{i})={\lambda_{i}}_{\ast}(\psi,\mathfrak{J})\in\mathfrak{A}(\alpha)$. Define the pair $(\varphi,\mathfrak{H})\in\mathfrak{A}(\alpha)$ by $\varphi|_{U_{i}\times V}=\psi^{i}|_{U_{i}\times V}$ and $\mathfrak{H}|_{U_{i}\times V}=\mathfrak{J}^{i}|_{U_{i}\times V}$ for every $1\leq i\leq n$. Then we have $d=\dim(\mathbf{H}_{u})$ for $\nu$-almost every $u\in U$. ∎
We then simply say that $(\varphi,\mathfrak{H})\in\mathfrak{A}(\alpha)$ is a minimal pair and we write $d=\dim(\mathbf{H}_{u})$ for $\nu$-almost every $u\in U$.
The following theorem is an extension of [Zi81, Lemma 2.1] to the case of algebraic groups defined over a local field $k$ of characteristic zero. For the sake of completeness, we give a proof following the one of [Zi81, Lemma 2.1] and we add a few details regarding the measurability of certain maps.
Theorem 3.3.
Keep the same notation as above. The following assertions hold:
  • $(\rm i)$
    Let $(\varphi,\mathfrak{H})\in\mathfrak{A}(\alpha)$ and $(\psi,\mathfrak{J})\in\mathfrak{A}(\alpha)$ be two minimal pairs. Then there exists a measurable map $U\to G:u\mapsto b_{u}$ such that $\mathbf{H}_{u}^{0}=b_{u}\mathbf{J}_{u}^{0}b_{u}^{-1}$ for $\nu$-almost every $u\in U$.
  • $(\rm ii)$
    Let $(\varphi,\mathfrak{H})\in\mathfrak{A}(\alpha)$ be a minimal pair. Then there exists a Zariski connected $k$-subgroup $\mathbf{H}<\mathbf{G}$ and a measurable map $U\to G:u\mapsto g_{u}$ such that $\mathbf{H}_{u}^{0}=g_{u}\mathbf{H}g_{u}^{-1}$ for $\nu$-almost every $u\in U$.
  • $(\rm iii)$
    Let $(\varphi,\mathfrak{H})\in\mathfrak{A}(\alpha)$ and $(\psi,\mathfrak{J})\in\mathfrak{A}(\alpha)$ be two minimal pairs for which there exists a Zariski connected $k$-subgroup $\mathbf{H}<\mathbf{G}$ such that $\mathbf{H}_{u}^{0}=\mathbf{H}=\mathbf{J}_{u}^{0}$ for $\nu$-almost every $u\in U$. Set $\mathbf{L}=\mathscr{N}_{\mathbf{G}}(\mathbf{H})$, $H=\mathbf{H}(k)$, $L=\mathbf{L}(k)$ and consider the factor map $p:G\to G/L$. Then we have $p\circ\varphi=p\circ\psi$ $\zeta$-almost everywhere.
Proof.
$(\rm i)$ After discarding a $\nu$-null Borel subset, we may assume that the Borel maps $\mathfrak{H}:U\to\operatorname{Sub}(G):u\mapsto H_{u}$ and $\mathfrak{J}:U\to\operatorname{Sub}(G):u\mapsto J_{u}$ satisfy that for every $u\in U$, there exist a $k$-subgroup $\mathbf{H}_{u}<\mathbf{G}$ such that $\mathbf{H}_{u}(k)=H_{u}$ and a $k$-subgroup $\mathbf{J}_{u}<\mathbf{G}$ such that $\mathbf{J}_{u}(k)=J_{u}$, and the Borel maps $V\to G/H_{u}:v\mapsto\varphi(u,v)H_{u}$ and $V\to G/J_{u}:v\mapsto\psi(u,v)J_{u}$ are $(\mathscr{V},\alpha_{u})$-equivariant. Denote by $\mathscr{G}=\operatorname{L}^{0}(V,\eta,G)$ the space of all $\eta$-equivalence classes of measurable functions $f:V\to G$. Endowed with the topology of convergence in measure, $\mathscr{G}$ is a Polish group. For every $u\in U$, consider the $k$-$\mathbf{G}$-variety $\mathbf{M}_{u}=\mathbf{G}/\mathbf{H}_{u}\times\mathbf{G}/\mathbf{J}_{u}$, where $\mathbf{G}\curvearrowright\mathbf{M}_{u}$ acts diagonally. Set $M_{u}=\mathbf{M}_{u}(k)$ and regard $N_{u}=G/H_{u}\times G/J_{u}\subset M_{u}$ as a closed subset. Define the Borel map
$\theta_{u}:V\to N_{u}:v\mapsto(\varphi(u,v)H_{u},\psi(u,v)J_{u}).$
Since $\mathbf{M}_{u}$ is an algebraic $k$-$\mathbf{G}$-variety, the continuous action $G\curvearrowright M_{u}$ has locally closed orbits (see Facts 2.5$(\rm i)$). Thus, the action $G\curvearrowright N_{u}$ has locally closed orbits and so the quotient space $G\backslash N_{u}$ is a standard Borel space. Since the measurable map
$V\to G\backslash N_{u}:v\mapsto G\cdot\theta_{u}(v)$
is $\mathscr{V}$-invariant and since $\mathscr{V}$ is ergodic, $V\to G\backslash N_{u}:v\mapsto G\cdot\theta_{u}(v)$ is $\eta$-almost everywhere constant. We may choose $b\in G$ such that $\theta_{u}(v)\in G\cdot(H_{u},bJ_{u})$ for $\eta$-almost every $v\in V$. Letting $L=\operatorname{Stab}_{G}(H_{u},bJ_{u})<G$, the map
$G/L\to G\cdot(H_{u},bJ_{u}):gL\mapsto g(H_{u},bJ_{u})$
is a homeomorphism. By considering a Borel section $G/L\to G$ of the projection map $G\to G/L$, we may then find $F\in\mathscr{G}$ such that $\theta_{u}(v)=(F(v)H_{u},F(v)bJ_{u})$ for $\eta$-almost every $v\in V$.
For every $u\in U$, define $\varphi_{u}:V\to G:v\mapsto\varphi(u,v)$, $\psi_{u}:V\to G:v\mapsto\psi(u,v)$ and regard $\varphi_{u}\in\mathscr{G}$, $\psi_{u}\in\mathscr{G}$. Then the maps $U\to\mathscr{G}:u\mapsto\varphi_{u}$ and $U\to\mathscr{G}:u\mapsto\psi_{u}$ are Borel. For every $u\in U$, set $\mathscr{H}_{u}=\operatorname{L}^{0}(V,\eta,H_{u})$ and $\mathscr{J}_{u}=\operatorname{L}^{0}(V,\eta,J_{u})$. By Proposition 2.9, the maps $U\to\operatorname{Sub}(\mathscr{G}):u\mapsto\mathscr{H}_{u}$ and $U\to\operatorname{Sub}(\mathscr{G}):u\mapsto\mathscr{J}_{u}$ are Borel. Denote by $\pi_{U}:U\times G\times\mathscr{G}\to U$ the projection map. Consider the subset
$B=\left\{(u,b,F)\in U\times G\times\mathscr{G}\mid F^{-1}\varphi_{u}\in\mathscr{H}_{u}\;\text{ and }\;b^{-1}F^{-1}\psi_{u}\in\mathscr{J}_{u}\right\}.$
The reasoning in the first paragraph shows that for every $u\in U$, there exist $F\in\mathscr{G}$ and $b\in G$ such that $(\varphi(u,v)H_{u},\psi(u,v)J_{u})=(F(v)H_{u},F(v)bJ_{u})$ for $\eta$-almost every $v\in V$. This means exactly that $(u,b,F)\in B$. Therefore, we have $\pi_{U}(B)=U$. Using (2.7), it follows that $B\subset U\times G\times\mathscr{G}$ is a Borel subset. By the measurable selection theorem (see Theorem 3.1), there exists a Borel map $U\to G\times\mathscr{G}:u\mapsto(b_{u},F_{u})$ such that $(u,b_{u},F_{u})\in B$ for $\nu$-almost every $u\in U$. Thus, we may choose Borel maps $\theta:W\to G$ and $\mathfrak{L}:U\mapsto\operatorname{Sub}(G):u\mapsto L_{u}$ such that $\theta(u,v)=F_{u}(v)$ for $\zeta$-almost every $(u,v)\in W$ and $L_{u}=H_{u}\cap b_{u}J_{u}b_{u}^{-1}$ for $\nu$-almost every $u\in U$. Note that for $\nu$-almost every $u\in U$, letting $\mathbf{L}_{u}=\mathbf{H}_{u}\cap b_{u}\mathbf{J}_{u}b_{u}^{-1}<\mathbf{G}$, which is a $k$-subgroup, we have $L_{u}=\mathbf{L}_{u}(k)$. For $\nu$-almost every $u\in U$, using the $G$-equivariant homeomorphism
$G/L_{u}\to G\cdot(H_{u},b_{u}J_{u}):gL_{u}\mapsto(gH_{u},gb_{u}J_{u}),$
we may identify
$\theta(u,v)L_{u}$ $=(\theta(u,v)H_{u},\theta(u,v)b_{u}J_{u})$
$=(F_{u}(v)H_{u},F_{u}(v)b_{u}J_{u})$
$=(\varphi(u,v)H_{u},\psi(u,v)J_{u}).$
for $\zeta$-almost every $(u,v)\in W$. Since for $\nu$-almost every $u\in U$, the Borel maps $V\to G/H_{u}:v\mapsto\varphi(u,v)H_{u}$ and $V\to G/J_{u}:v\mapsto\psi(u,v)J_{u}$ are $(\mathscr{V},\alpha_{u})$-equivariant, it follows that the Borel map $V\to G/L_{u}:v\mapsto\theta(u,v)L_{u}$ is $(\mathscr{V},\alpha_{u})$-equivariant. Therefore, we infer that $(\theta,\mathfrak{L})\in\mathfrak{A}(\alpha)$. By minimality of the pairs $(\varphi,\mathfrak{H})\in\mathfrak{A}(\alpha)$ and $(\psi,\mathfrak{J})\in\mathfrak{A}(\alpha)$, we have that $\dim(\mathbf{H}_{u})=\dim(\mathbf{H}_{u}\cap b_{u}\mathbf{J}_{u}b_{u}^{-1})=\dim(\mathbf{J}_{u})$ for $\nu$-almost every $u\in U$. This further implies that $\mathbf{H}_{u}^{0}=(\mathbf{H}_{u}\cap b_{u}\mathbf{J}_{u}b_{u}^{-1})^{0}=b_{u}\mathbf{J}_{u}^{0}b_{u}^{-1}$ for $\nu$-almost every $u\in U$.
$(\rm ii)$ Consider the adjoint $k$-representation $\operatorname{Ad}:\mathbf{G}\to\operatorname{GL}(\operatorname{Lie}(\mathbf{G}))$. Denote by $M=\operatorname{Gr}_{d}(\operatorname{Lie}(G))$ (resp. $\mathbf{M}=\operatorname{Gr}(\operatorname{Lie}(\mathbf{G}))$) the $k$-analytic Grassmannian $G$-manifold (resp. algebraic $k$-$\mathbf{G}$-variety) of all $d$-dimensional subspaces of $\operatorname{Lie}(G)$ (resp. $\operatorname{Lie}(\mathbf{G})$). We naturally have $M=\mathbf{M}(k)$ as $k$-analytic manifolds. The map $U\to M:u\mapsto\operatorname{Lie}(H_{u})$ is measurable (see [NZ00, Section 3]). Since $\mathbf{M}$ is an algebraic $k$-$\mathbf{G}$-variety, the continuous action $G\curvearrowright M$ has locally closed orbits and so the quotient space $G\backslash M$ is a standard Borel space (see Facts 2.5$(\rm i)$). Consider the measurable map $\beta:U\to G\backslash M:u\mapsto G\cdot\operatorname{Lie}(H_{u})$. Let $\lambda\in[\mathscr{U}]$. Since $\lambda_{\ast}(\varphi,\mathfrak{H})\in\mathfrak{A}(\alpha)$ is a minimal pair, item $(\rm i)$ implies that there exists a measurable map $U\to G:u\mapsto b_{u}$ such that $\mathbf{H}_{u}^{0}=b_{u}\mathbf{H}_{\lambda u}^{0}b_{u}^{-1}$ for $\nu$-almost every $u\in U$. Then we have $\beta(\lambda u)=\beta(u)$ for $\nu$-almost every $u\in U$. This implies that the measurable map $\beta:U\to G\backslash M$ is $\mathscr{U}$-invariant. Since $\mathscr{U}$ is ergodic, $\beta:U\to G\backslash M$ is $\nu$-almost everywhere constant and so there exists a point $w\in M$ such that $\operatorname{Lie}(H_{u})\in G\cdot w$ for $\nu$-almost every $u\in U$. Then there exists a measurable map $U\to G:u\mapsto g_{u}$ such that $\operatorname{Lie}(H_{u})=g_{u}\cdot w$ for $\nu$-almost every $u\in U$. By Fubini’s theorem, there exists $u_{0}\in U$ such that letting $\mathbf{H}=g_{u_{0}}^{-1}\mathbf{H}_{u_{0}}^{0}g_{u_{0}}$, we have $\operatorname{Lie}(\mathbf{H}_{u}^{0})=\operatorname{Lie}(g_{u}\mathbf{H}g_{u}^{-1})$ for $\nu$-almost every $u\in U$. Since the characteristic of $k$ is zero, it follows that $\mathbf{H}_{u}^{0}=g_{u}\mathbf{H}g_{u}^{-1}$ for $\nu$-almost every $u\in U$.
$(\rm iii)$ By the proof of item $(\rm i)$, there exist measurable maps $\theta:W\to G$ and $U\to G:u\mapsto b_{u}$ such that
$(\varphi(u,v)H_{u},\psi(u,v)J_{u})=(\theta(u,v)H_{u},\theta(u,v)b_{u}J_{u})$
for $\zeta$-almost every $(u,v)\in W$. For $\nu$-almost every $u\in U$, since $\dim(\mathbf{H}_{u})=\dim(\mathbf{H}_{u}\cap b_{u}\mathbf{J}_{u}b_{u}^{-1})=\dim(\mathbf{J}_{u})$ and $\mathbf{H}_{u}^{0}=\mathbf{H}=\mathbf{J}_{u}^{0}$, we have $\mathbf{H}=\mathbf{H}_{u}^{0}=b_{u}\mathbf{J}_{u}^{0}b_{u}^{-1}=b_{u}\mathbf{H}b_{u}^{-1}$ and so $b_{u}\in\mathscr{N}_{\mathbf{G}}(\mathbf{H})=\mathbf{L}$. Then for $\zeta$-almost every $(u,v)\in W$, we have
$\varphi(u,v)L=\theta(u,v)L=\theta(u,v)b_{u}L=\psi(u,v)L.$
This shows that $p\circ\varphi=p\circ\psi$ $\zeta$-almost everywhere. ∎
3.2. Proof of Theorem A
Let $k$ be a local field of characteristic zero, $\mathbf{G}$ a $k$-isotropic almost $k$-simple algebraic $k$-group and $\Gamma<\mathbf{G}(k)$ a Zariski dense discrete subgroup. Let $\Gamma\curvearrowright(X,\nu)$ be a free ergodic pmp action and set $\mathscr{R}=\mathscr{R}(\Gamma\curvearrowright X)$. Upon discarding a null Borel subset, we may assume that the Borel action $\Gamma\curvearrowright X$ is free. Consider the orbit Borel $1$-cocycle $\beta:\mathscr{R}\to\Gamma:(\gamma x,x)\mapsto\gamma$. Let $\mathscr{S}$ be an ergodic equivalence relation on $(Y,\eta)$ and consider the Borel $1$-cocycle $\alpha=\beta\circ q:\mathscr{R}\times\mathscr{S}\to\Gamma$, where $q:\mathscr{R}\times\mathscr{S}\to\mathscr{R}$ is the canonical factor map. Then $\ker(\alpha)=\operatorname{id}_{X}\times\mathscr{S}$.
For every $i\in\{1,2\}$, let $\mathscr{T}_{i}$ be an ergodic equivalence relation on $(Z_{i},\zeta_{i})$. Set $\mathscr{T}=\mathscr{T}_{1}\times\mathscr{T}_{2}$ and $(Z,\zeta)=(Z_{1}\times Z_{2},\zeta_{1}\otimes\zeta_{2})$. Assume that ×$\mathscr{R}\times\mathscr{S}\cong\mathscr{T}$. Then there exists a conull Borel subset $Z_{0}\subset Z$ such that we may identify (×$(\mathscr{R}\times\mathscr{S})|_{Z_{0}}$ with $\mathscr{T}|_{Z_{0}}$ everywhere and regard $Z_{0}\subset X\times Y$ as a conull Borel subset.
We prove Theorem A by contradiction. Assume that for every $i\in\{1,2\}$, we have $\mathscr{T}_{i}\npreceq_{\mathscr{T}}\mathscr{S}$ as subequivalences relations of $\mathscr{T}$.
By [Zi84, Proposition 9.3.2], for every $i\in\{1,2\}$, there exists an amenable ergodic subequivalence relation $\mathscr{U}_{i}\leq\mathscr{T}_{i}$. Set $\mathscr{U}=\mathscr{U}_{1}\times\mathscr{U}_{2}$ and $\mathscr{U}_{0}=\mathscr{U}|_{Z_{0}}$ which is an amenable ergodic subequivalence relation of $\mathscr{T}|_{Z_{0}}$.
Claim 3.4.
There exist a proper $k$-subgroup $\mathbf{H}<\mathbf{G}$ and a $(\mathscr{U},\alpha|_{\mathscr{U}})$-equivariant measurable map $f:Z\to G/H$, where we regard $G/H\subset(\mathbf{G}/\mathbf{H})(k)$ with $H=\mathbf{H}(k)$.
By contradiction, assume that the assertion on Claim 3.4 does not hold. By applying Theorem 2.6 to $\alpha|_{\mathscr{U}_{0}}:\mathscr{U}_{0}\to G$, there exists a compact subgroup $L<G$ and a $(\mathscr{U}_{0},\alpha|_{\mathscr{U}_{0}})$-equivariant measurable map $f:Z_{0}\to G/L$. Upon discarding a null Borel subset, we may assume that the map $f:Z_{0}\to G/L$ is Borel and strictly $(\mathscr{U}_{0},\alpha|_{\mathscr{U}_{0}})$-equivariant. Choose a Borel section $\sigma:G/L\to G$ and define the Borel map $\psi=\sigma\circ f:Z_{0}\to G$. Then for every $(z_{1},z_{2})\in\mathscr{U}|_{Z_{0}}$, we have $\psi(z_{1})L=\alpha(z_{1},z_{2})\psi(z_{2})L$. Since $G$ is $\sigma$-compact, there exists a compact subset $C\subset G$ such that the Borel subset $V=\psi^{-1}(C)\subset Z_{0}$ satisfies $\zeta(V)>0$. Then for every $(z_{1},z_{2})\in\mathscr{U}|_{V}$, we have $\alpha(z_{1},z_{2})\in\psi(z_{1})L\psi(z_{2})^{-1}\subset CLC^{-1}\cap\Gamma$. Since $CLC^{-1}\subset G$ is compact and $\Gamma<G$ is discrete, $F=CLC^{-1}\cap\Gamma$ is a finite subset of $\Gamma$.
Define the subequivalence relation $\mathscr{V}=\mathscr{U}|_{V}\cap\ker(\alpha)|_{V}$. Let $z\in V$. Then we can write $[z]_{\mathscr{U}|_{V}}=\bigcup_{\gamma\in F}C_{\gamma}(z)$, where $C_{\gamma}(z)=\left\{t\in[z]_{\mathscr{U}|_{V}}\mid\alpha(z,t)=\gamma\right\}$ for $\gamma\in F$. Then for every $\gamma\in F$ and all $t_{1},t_{2}\in C_{\gamma}(z)$, we have $\alpha(t_{1},t_{2})=\alpha(t_{1},z)\alpha(z,t_{2})=\gamma^{-1}\gamma=e$ and so $(t_{1},t_{2})\in\mathscr{V}$. Therefore, $\mathscr{V}\subset\mathscr{U}|_{V}$ has bounded index. Since $\mathscr{V}\leq\ker(\alpha)|_{V}=(\operatorname{id}_{X}\times\mathscr{S})|_{V}$, it follows that $\mathscr{U}\preceq_{\mathscr{T}}\operatorname{id}_{X}\times\mathscr{S}$. Since $\mathscr{U}_{1}\times\operatorname{id}_{Z_{2}}\leq\mathscr{U}$, this further implies that $\mathscr{U}_{1}\times\operatorname{id}_{Z_{2}}\preceq_{\mathscr{T}}\operatorname{id}_{X}\times\mathscr{S}$. Since $\mathscr{U}_{1}$ is ergodic, Facts 2.2$(\rm v)$ implies that $\mathscr{T}_{1}\times\operatorname{id}_{Z_{2}}\preceq_{\mathscr{T}}\operatorname{id}_{X}\times\mathscr{S}$. This contradicts our assumption and finishes the proof of Claim 3.4.
Claim 3.5.
There exist a proper $k$-subgroup $\mathbf{L}<\mathbf{G}$ and a $(\mathscr{T}_{1}\times\mathscr{T}_{2},\alpha)$-equivariant measurable map $F:Z\to G/L$, where we regard $G/L\subset(\mathbf{G}/\mathbf{L})(k)$ with $L=\mathbf{L}(k)$.
The proof of Claim 3.5 consists of two steps. Firstly, we consider the product equivalence $\mathscr{T}_{1}\times\mathscr{U}_{2}$ together with the subequivalence relation $\mathscr{U}_{1}\times\mathscr{U}_{2}\leq\mathscr{T}_{1}\times\mathscr{U}_{2}$ and the $(\mathscr{U}_{1}\times\mathscr{U}_{2},\alpha|_{\mathscr{U}_{1}\times\mathscr{U}_{2}})$-equivariant measurable map $f:Z\to G/H$. By Theorem 3.3$(\rm ii)$, we may choose a minimal pair $(\psi,\mathfrak{J})\in\mathfrak{A}(\alpha|_{\mathscr{T}_{1}\times\mathscr{U}_{2}})$ for which there exists a Zariski connected $k$-subgroup $\mathbf{J}<\mathbf{G}$ such that $\mathbf{J}=\mathbf{J}_{z_{1}}^{0}$ for $\zeta_{1}$-almost every $z_{1}\in Z_{1}$. By Claim 3.4, there exist a proper $k$-subgroup $\mathbf{H}<\mathbf{G}$ and a $(\mathscr{U},\alpha|_{\mathscr{U}})$-equivariant measurable map $f:Z\to G/H$. Choose a Borel section $\sigma:G/H\to G$. Consider the measurable map $\varphi=\sigma\circ f:Z\to G$ and the constant map $\mathfrak{H}:Z_{1}\to\operatorname{Sub}(G):z_{1}\mapsto H$. Then $(\varphi,\mathfrak{H})\in\mathfrak{A}(\alpha|_{\mathscr{T}_{1}\times\mathscr{U}_{2}})$. By minimality of $(\psi,\mathfrak{J})\in\mathfrak{A}(\alpha|_{\mathscr{T}_{1}\times\mathscr{U}_{2}})$ and since $\mathbf{H}<\mathbf{G}$ is a proper $k$-subgroup, it follows that $\mathbf{J}<\mathbf{G}$ is a proper $k$-subgroup.
Next, we claim that $\mathbf{J}\neq\{e\}$. Upon discarding a null Borel subset, we may assume that the map $\mathfrak{J}:Z_{1}\to\operatorname{Sub}(G):z_{1}\mapsto J_{z_{1}}$ is Borel and for every $z_{1}\in Z_{1}$, there exists an algebraic $k$-subgroup $\mathbf{J}_{z_{1}}<\mathbf{G}$ such that $\mathbf{J}_{z_{1}}(k)=J_{z_{1}}$ and $\mathbf{J}_{z_{1}}^{0}=\mathbf{J}$. By contradiction, assume that $\mathbf{J}=\{e\}$. Then we have that $J_{z_{1}}<G$ is a finite subgroup for every $z_{1}\in Z_{1}$. It is well-known that there are only finitely many conjugacy classes of maximal compact subgroups of $G$ (see e.g. [BT71, Corollaire 3.3.3]). Therefore, there exists a nonnull Borel subset $W\subset Z_{1}$ and a maximal compact subgroup $K<G$ such that the subset
$B=\left\{(z_{1},b)\in W\times G\mid bJ_{z_{1}}b^{-1}<K\right\}\subset W\times G$
satisfies $\pi_{W}(B)=W$. Using (2.6), it follows that $B\subset W\times G$ is a Borel subset. By the measurable selection theorem (see Theorem 3.1), there exists a Borel map $W\to G:z_{1}\mapsto b_{z_{1}}$ such that $b_{z_{1}}J_{z_{1}}b_{z_{1}}^{-1}<K$ for $\zeta_{1}$-almost every $z_{1}\in W$. Upon discarding a null Borel subset, we may assume that the Borel map $W\to G:z_{1}\mapsto b_{z_{1}}$ satisfies $b_{z_{1}}J_{z_{1}}b_{z_{1}}^{-1}<K$ for every $z_{1}\in W$. We may extend the Borel map $Z_{1}\to G:z_{1}\mapsto b_{z_{1}}$ by declaring that $b_{z_{1}}=e$ for every $z_{1}\in Z_{1}\setminus W$. Upon replacing the pair $(\psi,\mathfrak{J})\in\mathfrak{A}(\alpha|_{\mathscr{T}_{1}\times\mathscr{U}_{2}})$ by the pair $(\rho,\mathfrak{K})\in\mathfrak{A}(\alpha|_{\mathscr{T}_{1}\times\mathscr{U}_{2}})$, where $\rho:Z\to G:(z_{1},z_{2})\mapsto\psi(z_{1},z_{2})b_{z_{1}}^{-1}$ and $\mathfrak{K}:Z_{1}\to\operatorname{Sub}(G):z_{1}\mapsto b_{z_{1}}J_{z_{1}}b_{z_{1}}^{-1}$, we may assume that $J_{z_{1}}<K$ for every $z_{1}\in W$.
Choose a conull Borel subset $V\subset(W\times Z_{2})\cap Z_{0}$ such that the restricted Borel map $\psi|_{V}:V\to G$ satisfies
$\forall((z_{1},s),(z_{1},t))\in(\operatorname{id}_{Z_{1}}\times\mathscr{U}_{2})|_{V},\quad\psi(z_{1},s)J_{z_{1}}=\alpha((z_{1},s),(z_{1},t))\psi(z_{1},t)J_{z_{1}}.$
Since $G$ is $\sigma$-compact, there exists a compact subset $C\subset G$ such that $V_{0}=(\psi|_{V})^{-1}(C)\subset V$ satisfies $\zeta(V_{0})>0$. Then for every $((z_{1},s),(z_{1},t))\in(\operatorname{id}_{Z_{1}}\times\mathscr{U}_{2})|_{V_{0}}$, we have
$\alpha((z_{1},s),(z_{1},t))\in\psi(z_{1},s)J_{z_{1}}\psi(z_{1},t)^{-1}\in CJ_{z_{1}}C^{-1}\cap\Gamma\subset CKC^{-1}\cap\Gamma.$
Since $CKC^{-1}\subset G$ is compact and $\Gamma<G$ is discrete, $F=CKC^{-1}\cap\Gamma$ is a finite subset of $\Gamma$. Now a reasoning entirely analogous to the one as in the second paragraph of the proof of Claim 3.4 shows that $\operatorname{id}_{Z_{1}}\times\mathscr{U}_{2}\preceq_{\mathscr{T}}\operatorname{id}_{X}\times\mathscr{S}$. This contradicts our assumption and so $\mathbf{J}\neq\{e\}$.
Since $\mathbf{G}$ is almost $k$-simple and since the Zariski connected $k$-subgroup $\mathbf{J}$ is different from $\{e\}$ and $\mathbf{G}$, it follows that the $k$-subgroup $\mathbf{L}=\mathscr{N}_{\mathbf{G}}(\mathbf{J})<\mathbf{G}$ is a proper subgroup. Set $L=\mathbf{L}(k)$. Choose a countable subgroup $\Lambda<[\mathscr{T}_{1}]$ such that $\mathscr{T}_{1}=\mathscr{R}(\Lambda\curvearrowright Z_{1})$. For every $\lambda\in\Lambda$, the pair $\lambda_{\ast}(\psi,\mathfrak{J})\in\mathfrak{A}(\alpha|_{\mathscr{T}_{1}\times\mathscr{U}_{2}})$ is minimal and for $\zeta_{1}$-almost every $z_{1}\in Z_{1}$, we have $(\mathbf{J}_{\lambda z_{1}})^{0}=\mathbf{J}$. Then Theorem 3.3$(\rm iii)$ implies that for $\zeta$-almost every $(z_{1},s)\in Z_{1}\times Z_{2}$, we have $\psi(z_{1},s)L=\alpha((z_{1},s),(\lambda z_{1},s))\psi(\lambda z_{1},s)L$. This further implies that for almost every $(z_{1},s_{1})\in\mathscr{T}_{1}$ and almost every $(z_{2},s_{2})\in\mathscr{U}_{2}$, we have
$\psi(z_{1},s_{1})L$ $=\alpha((z_{1},s_{1}),(z_{2},s_{1}))\psi(z_{2},s_{1})L$
$=\alpha((z_{1},s_{1}),(z_{2},s_{1}))\alpha((z_{2},s_{1}),(z_{2},s_{2}))\psi(z_{2},s_{2})L$
$=\alpha((z_{1},s_{1}),(z_{2},s_{2}))\psi(z_{2},s_{2})L,$
where in the second line we used the fact that $(\psi,\mathfrak{J})\in\mathfrak{A}(\alpha|_{\mathscr{T}_{1}\times\mathscr{U}_{2}})$. It follows that the measurable map $Z\to G/L:z\mapsto\psi(z)L$ is $(\mathscr{T}_{1}\times\mathscr{U}_{2},\alpha|_{\mathscr{T}_{1}\times\mathscr{U}_{2}})$-equivariant.
Now, we can consider the product equivalence relation $\mathscr{T}_{1}\times\mathscr{T}_{2}$ together with the subequivalence relation $\mathscr{T}_{1}\times\mathscr{U}_{2}\leq\mathscr{T}_{1}\times\mathscr{T}_{2}$ and the $(\mathscr{T}_{1}\times\mathscr{U}_{2})$-equivariant measurable map $Z\to G/L:z\mapsto\psi(z)L$. Repeating the same argument as in the first four paragraphs of the proof of Claim 3.5 and since $\mathscr{T}_{1}\times\operatorname{id}_{Z_{2}}\npreceq_{\mathscr{T}}\operatorname{id}_{X}\times\mathscr{S}$, upon modifying the proper $k$-subgroup $\mathbf{L}<\mathbf{G}$, we obtain the desired conclusion. This finishes the proof of Claim 3.5.
Consider now the pmp action $\Gamma\curvearrowright(Z,\zeta)$ defined by $\gamma z=(\gamma x,y)$ for every $z=(x,y)\in Z$. Then for every $\gamma\in\Gamma$ and almost every $z=(x,y)\in Z$, we have
$F(\gamma z)=F(\gamma x,y)=\alpha((\gamma x,y),(x,y))F(x,y)=\gamma F(z).$
Then the measurable map $F:Z\to G/L$ is $\Gamma$-equivariant. Consider the Borel probability measure $\mu=F_{\ast}\zeta\in\operatorname{Prob}(G/L)^{\Gamma}$. Since $\Gamma<G$ is Zariski dense in $\mathbf{G}$, [BDL14, Proposition 1.9] implies that there exists a normal $k$-subgroup $\mathbf{N}\lhd\mathbf{G}$ such that the image of $\Gamma$ in $(\mathbf{G}/\mathbf{N})(k)$ is precompact and $\mu$ is supported on $(\mathbf{G}/\mathbf{L})^{\mathbf{N}}\cap G/L$. Since $\mathbf{L}<\mathbf{G}$ is a proper $k$-subgroup and since $(\mathbf{G}/\mathbf{L})^{\mathbf{N}}\neq\emptyset$, we have $\mathbf{N}\neq\mathbf{G}$. Since $\mathbf{G}$ is almost $k$-simple, it follows that $\mathbf{N}\lhd\mathbf{G}$ is a finite normal $k$-subgroup. Letting $N=\mathbf{N}(k)$, the image of $\Gamma$ in $(\mathbf{G}/\mathbf{N})(k)$ is contained in $G/N$ which is closed in $(\mathbf{G}/\mathbf{N})(k)$. Then the image of $\Gamma$ in $G/N$ is precompact. Since $N<G$ is finite, this further implies that $\Gamma<G$ is compact, which is a contradiction. Therefore, we have showed that the orbit equivalence relation $\mathscr{R}(\Gamma\curvearrowright X)$ is strongly prime. This concludes the proof of Theorem A.
The following variation of Theorem A will be useful in the proof of Theorem D.
Theorem 3.6.
Let $k$ be a local field of characteristic zero, $\mathbf{G}$ a $k$-isotropic almost $k$-simple algebraic $k$-group and $\Gamma<\mathbf{G}(k)$ a Zariski dense discrete subgroup. Let $\Lambda$ be a countable discrete group and $\Gamma\times\Lambda\curvearrowright(X,\nu)$ an essentially free ergodic pmp action. Set $\mathscr{R}=\mathscr{R}(\Gamma\times\Lambda\curvearrowright X)$.
For every $i\in\{1,2\}$, let $\mathscr{T}_{i}$ be an ergodic equivalence relation on $(Z_{i},\zeta_{i})$ and assume that $\mathscr{R}\cong\mathscr{T}_{1}\times\mathscr{T}_{2}$. Then there exists $i\in\{1,2\}$ such that $\mathscr{T}_{i}\preceq_{\mathscr{R}}\mathscr{R}(\Lambda\curvearrowright X)$ as subequivalence relations.
Proof.
The proof is identical to the proof of Theorem A, the only difference is that the pmp action $\Gamma\times\Lambda\curvearrowright(X,\nu)$ is no longer assumed to be a product action. To circumvent this issue, we exploit Lemma 2.3. Upon discarding a null Borel subset, we may assume that the Borel action $\Gamma\times\Lambda\curvearrowright X$ is free. Consider the Borel $1$-cocycle $\alpha:\mathscr{R}\to\Gamma:((\gamma,\lambda)x,x)\mapsto\gamma$. Then $\ker(\alpha)=\mathscr{R}(\Lambda\curvearrowright X)$.
By contradiction, assume that for every $i\in\{1,2\}$, we have $\mathscr{T}_{i}\npreceq_{\mathscr{R}}\mathscr{R}(\Lambda\curvearrowright X)$ as subequivalence relations. By Lemma 2.3, for every $i\in\{1,2\}$, there exists an amenable ergodic subequivalence relation $\mathscr{U}_{i}\leq\mathscr{T}_{i}$ such that $\mathscr{U}_{i}\npreceq_{\mathscr{R}}\mathscr{R}(\Lambda\curvearrowright X)$. Then we can literally repeat the exact same proof as in Claims 3.4 and 3.5 to deduce a contradiction. ∎
We use Theorem 3.6 in combination with Connes–Jones’ construction [CJ81] to provide examples of type ${\rm II_{1}}$ ergodic equivalence relations $\mathscr{R}$ that are prime but for which the associated type ${\rm II_{1}}$ factor $\operatorname{L}(\mathscr{R})$ is not prime.
Proposition 3.7.
Let $\Gamma=\operatorname{SL}_{3}(\mathbb{Z})$ and $H$ be a nonabelian finite discrete group. Set $\Lambda=\bigoplus_{\mathbb{N}}H$, $(Y,\eta)=(H^{\mathbb{N}},\operatorname{Haar})$ and $(X,\nu)=(Y^{\Gamma},\eta^{\otimes\Gamma})$. Consider the free ergodic pmp action $\Gamma\times\Lambda\curvearrowright(X,\nu)$, where $\Lambda\curvearrowright(X,\nu)$ acts diagonally and $\Gamma\curvearrowright(X,\nu)$ acts by Bernoulli shifts. Set $\mathscr{R}=\mathscr{R}(\Gamma\times\Lambda\curvearrowright X)$.
Then $\mathscr{R}$ is prime but $\operatorname{L}(\mathscr{R})$ is not prime. In fact, $\operatorname{L}(\mathscr{R})$ is a McDuff factor meaning that $\operatorname{L}(\mathscr{R})\cong\operatorname{L}(\mathscr{R})\mathbin{\overline{\otimes}}R$, where $R$ is the unique hyperfinite type ${\rm II_{1}}$ factor.
Proof.
By Connes–Jones’ result [CJ81], $\operatorname{L}(\mathscr{R})$ is a McDuff factor. Since $\Gamma$ has Kazhdan’s property (T) and since $\Gamma\curvearrowright(X,\nu)$ is ergodic, it follows that $\mathscr{R}$ is strongly ergodic. Assume that $\mathscr{R}\cong\mathscr{T}_{1}\times\mathscr{T}_{2}$, where $\mathscr{T}_{1}$, $\mathscr{T}_{2}$ are ergodic equivalence relations. By Theorem 3.6, there exists $i\in\{1,2\}$ such that $\mathscr{T}_{i}\preceq_{\mathscr{R}}\mathscr{R}(\Lambda\curvearrowright X)$ as subequivalence relations. Since $\mathscr{R}(\Lambda\curvearrowright X)$ is amenable, it follows that $\mathscr{T}_{i}$ is amenable and so $\mathscr{T}_{i}$ is hyperfinite by [CFW81]. Since $\mathscr{R}$ is strongly ergodic, it follows that $\mathscr{T}_{i}$ is finite. Therefore, $\mathscr{R}$ is prime. ∎
3.3. Proof of Corollary B
Let $n\geq 1$ and $\mathscr{R}_{1},\dots,\mathscr{R}_{n}$ be ergodic equivalence relations that belong to $\mathfrak{Z}$. For every $i\in\{1,\dots,n\}$, write $\mathscr{R}_{i}=\mathscr{R}(\Gamma_{i}\curvearrowright X_{i})$.
$(\rm i)$ Assume that $\mathscr{R}=\mathscr{R}_{1}\times\cdots\times\mathscr{R}_{n}\cong\mathscr{T}_{1}\times\mathscr{T}_{2}$, where $\mathscr{T}_{1}$ and $\mathscr{T}_{2}$ are type ${\rm II_{1}}$ ergodic equivalence relations acting on $(Y_{1},\eta_{1})$ and $(Y_{2},\eta_{2})$ respectively. For every $j\in\{1,2\}$, denote by $T_{j}\subset\{1,\dots,n\}$ a minimal subset for which $\mathscr{T}_{j}\preceq_{\mathscr{R}}\mathscr{R}_{T_{j}}$ as subequivalence relations, where $\mathscr{R}_{T_{j}}=\prod_{i\in T_{j}}\mathscr{R}_{i}$. Since $\mathscr{T}_{j}$ is a type ${\rm II_{1}}$ ergodic equivalence relation, we have $T_{j}\neq\emptyset$. Set $T=T_{1}\cup T_{2}$ and $\mathscr{R}_{T}=\prod_{i\in T}\mathscr{R}_{i}$.
Firstly, we claim that $T=\{1,\dots,n\}$. Indeed, we have $\mathscr{T}_{1}\preceq_{\mathscr{R}}\mathscr{R}_{T}$ and $\mathscr{T}_{2}\preceq_{\mathscr{R}}\mathscr{R}_{T}$ as subequivalence relations. By Facts 2.2$(\rm i)$, we obtain that $\operatorname{L}(\mathscr{T}_{1})\mathbin{\overline{\otimes}}\operatorname{L}^{\infty}(Y_{2})\preceq_{\operatorname{L}(\mathscr{R})}\operatorname{L}(\mathscr{R}_{T})\mathbin{\overline{\otimes}}\operatorname{L}^{\infty}(X_{T^{c}})$ and $\operatorname{L}^{\infty}(Y_{1})\mathbin{\overline{\otimes}}\operatorname{L}(\mathscr{T}_{2})\preceq_{\operatorname{L}(\mathscr{R})}\operatorname{L}(\mathscr{R}_{T})\mathbin{\overline{\otimes}}\operatorname{L}^{\infty}(X_{T^{c}})$. By [DHI16, Lemma 2.4(3)], we moreover have $\operatorname{L}(\mathscr{T}_{1})\mathbin{\overline{\otimes}}\operatorname{L}^{\infty}(Y_{2})\preceq^{s}_{\operatorname{L}(\mathscr{R})}\operatorname{L}(\mathscr{R}_{T})\mathbin{\overline{\otimes}}\operatorname{L}^{\infty}(X_{T^{c}})$ and $\operatorname{L}^{\infty}(Y_{1})\mathbin{\overline{\otimes}}\operatorname{L}(\mathscr{T}_{2})\preceq^{s}_{\operatorname{L}(\mathscr{R})}\operatorname{L}(\mathscr{R}_{T})\mathbin{\overline{\otimes}}\operatorname{L}^{\infty}(X_{T^{c}})$. Then [BV12, Lemma 2.3] implies that $\operatorname{L}(\mathscr{R})\preceq_{\operatorname{L}(\mathscr{R})}\operatorname{L}(\mathscr{R}_{T})\mathbin{\overline{\otimes}}\operatorname{L}^{\infty}(X_{T^{c}})$. Since the group $\Gamma_{i}$ is infinite and the pmp action $\Gamma_{i}\curvearrowright(X_{i},\nu_{i})$ is essentially free ergodic for every $i\in\{1,\dots,n\}$, [HI17, Theorem 4.4] implies that we necessarily have $T=\{1,\dots,n\}$.
Secondly, we claim that $T_{1}\cap T_{2}=\emptyset$. By contradiction, assume that $T_{1}\cap T_{2}\neq\emptyset$ and let $i\in T_{1}\cap T_{2}$. Write $\mathscr{R}=\mathscr{R}_{i}\times\mathscr{S}$, where $\mathscr{S}=\mathscr{R}_{\{1,\dots,n\}\setminus\{i\}}$. By Theorem A, we have $\mathscr{T}_{1}\preceq_{\mathscr{R}}\mathscr{S}$ or $\mathscr{T}_{2}\preceq_{\mathscr{R}}\mathscr{S}$ as subequivalence relations. Assume that $\mathscr{T}_{1}\preceq_{\mathscr{R}}\mathscr{S}$ as subequivalence relations. The reasoning in the previous paragraph shows that $\operatorname{L}(\mathscr{T}_{1})\mathbin{\overline{\otimes}}\operatorname{L}^{\infty}(Y_{2})\preceq^{s}_{\operatorname{L}(\mathscr{R})}\operatorname{L}(\mathscr{S})\mathbin{\overline{\otimes}}\operatorname{L}^{\infty}(X_{i})$ and $\operatorname{L}(\mathscr{T}_{1})\mathbin{\overline{\otimes}}\operatorname{L}^{\infty}(Y_{2})\preceq^{s}_{\operatorname{L}(\mathscr{R})}\operatorname{L}(\mathscr{R}_{T_{1}})\mathbin{\overline{\otimes}}\operatorname{L}^{\infty}(X_{T_{1}^{c}})$. Then [DHI16, Lemma 2.8(2)] implies that $\operatorname{L}(\mathscr{T}_{1})\mathbin{\overline{\otimes}}\operatorname{L}^{\infty}(Y_{2})\preceq^{s}_{\operatorname{L}(\mathscr{R})}\operatorname{L}(\mathscr{R}_{T_{1}\setminus\{i\}})\mathbin{\overline{\otimes}}\operatorname{L}^{\infty}(X_{(T_{1}\setminus\{i\})^{c}})$. Then Fact 2.2$(\rm i)$ implies that $\mathscr{T}_{1}\preceq_{\mathscr{R}}\mathscr{R}_{T_{1}\setminus\{i\}}$ as subequivalence relations. This however contradicts the minimality of the set $T_{1}$. Likewise, $\mathscr{T}_{2}\preceq_{\mathscr{R}}\mathscr{S}$ contradicts the minimality of the set $T_{2}$. Therefore, we have $T_{1}\cap T_{2}=\emptyset$.
Therefore, we may write $\{1,\dots,n\}=T_{1}\sqcup T_{2}$. We have $\mathscr{R}=\mathscr{R}_{T_{1}}\times\mathscr{R}_{T_{2}}\cong\mathscr{T}_{1}\times\mathscr{T}_{2}$. Moreover, we have $\mathscr{T}_{1}\preceq_{\mathscr{R}}\mathscr{R}_{T_{1}}$ and $\mathscr{T}_{2}\preceq_{\mathscr{R}}\mathscr{R}_{T_{2}}$ as subequivalence relations.
$(\rm ii)$ We prove the assertion by complete induction over $n\geq 1$. Theorem A implies that the assertion holds for $n=1$. Assume that the assertion holds for every $11\leq j\leq n-$ 1 and let us prove that it holds for $n$.
Firstly, assume that $\mathscr{R}=\mathscr{R}_{1}\times\cdots\times\mathscr{R}_{n}\cong\mathscr{T}_{1}\times\cdots\times\mathscr{T}_{p}$, where $\mathscr{T}_{1},\dots,\mathscr{T}_{p}$ are type ${\rm II_{1}}$ ergodic equivalence relations. We may assume that $p\geq n\geq 2$. Then item $(\rm i)$ implies that there exists a partition $\{1,\dots,n\}=T_{1}\sqcup T_{2}$ into nonempty subsets such that $\mathscr{T}_{1}\preceq_{\mathscr{R}}\mathscr{R}_{T_{1}}$ and $\mathscr{T}_{2}\times\cdots\times\mathscr{T}_{p}\preceq_{\mathscr{R}}\mathscr{R}_{T_{2}}$ as subequivalence relations. By Facts 2.2$(\rm vi)$, there exists an ergodic equivalence relation $\mathscr{V}$ such that $\mathscr{T}_{2}\times\cdots\times\mathscr{T}_{p}\times\mathscr{V}$ and $\mathscr{R}_{T_{2}}$ are stably isomorphic. Upon replacing $\mathscr{T}_{2}$ by $\mathscr{T}_{2}^{t}$ for $t>0$, we may assume that $\mathscr{T}_{2}\times\cdots\times\mathscr{T}_{p}\times\mathscr{V}\cong\mathscr{R}_{T_{2}}$. By induction hypothesis and since $p-1\geq|T_{2}|$ and $|T_{2}|\leq n-1$, we have that $\mathscr{V}$ is finite, $p-1=|T_{2}|\leq n-1$ and so $p=n$.
Secondly, assume that $\mathscr{R}=\mathscr{R}_{1}\times\cdots\times\mathscr{R}_{n}\cong\mathscr{T}_{1}\times\cdots\times\mathscr{T}_{n}$. Then item $(\rm i)$ implies that there exists a partition $\{1,\dots,n\}=T_{1}\sqcup T_{2}$ into nonempty subsets such that $\mathscr{T}_{1}\preceq_{\mathscr{R}}\mathscr{R}_{T_{1}}$ and $\mathscr{T}_{2}\times\cdots\times\mathscr{T}_{n}\preceq_{\mathscr{R}}\mathscr{R}_{T_{2}}$ as subequivalence relations. By Facts 2.2$(\rm vi)$, there exists an ergodic equivalence relation $\mathscr{U}$ such that $\mathscr{T}_{1}\times\mathscr{U}$ and $\mathscr{R}_{T_{1}}$ are stably isomorphic and there exists an ergodic equivalence relation $\mathscr{V}$ such that $\mathscr{T}_{2}\times\cdots\times\mathscr{T}_{n}\times\mathscr{V}$ and $\mathscr{R}_{T_{2}}$ are stably isomorphic. Upon replacing $\mathscr{T}_{2}$ by $\mathscr{T}_{2}^{t}$ for $t>0$, we may assume that $\mathscr{T}_{2}\times\cdots\times\mathscr{T}_{n}\times\mathscr{V}\cong\mathscr{R}_{T_{2}}$. Since $|T_{2}|\leq n-1$, the previous paragraph implies that $\mathscr{V}$ is finite, $|T_{2}|=n-1$ and so $|T_{1}|=1$. Upon replacing $\mathscr{T}_{n}$ by $\mathscr{T}_{n}\times\mathscr{V}$, we may assume that $\mathscr{V}$ is trivial. Upon permuting the indices, we may assume that $T_{1}=\{1\}$ and so $\mathscr{T}_{1}\times\mathscr{U}$ and $\mathscr{R}_{1}$ are stably isomorphic. Since $\mathscr{R}_{1}$ is strongly prime, we have that $\mathscr{U}$ is finite and so $\mathscr{T}_{1}$ and $\mathscr{R}_{1}$ are stably isomorphic. Moreover, we have $\mathscr{R}_{2}\times\cdots\times\mathscr{R}_{n}\cong\mathscr{T}_{2}\times\cdots\times\mathscr{T}_{n}$. By induction hypothesis and upon permuting the indices, we have that $\mathscr{T}_{i}$ and $\mathscr{R}_{i}$ are stably isomorphic for every $i\in\{2,\dots,n\}$. This concludes the proof of Corollary B.
3.4. Proof of Corollary C
Let $n\geq 1$ and $\mathscr{R}_{1},\dots,\mathscr{R}_{n}$ be strongly ergodic equivalence relations that belong to $\mathfrak{Z}$. For every $i\in\{1,\dots,n\}$, write $\mathscr{R}_{i}=\mathscr{R}(\Gamma_{i}\curvearrowright X_{i})$. For every $i\in\{1,\dots,n\}$, since $\Gamma_{i}$ is not inner amenable by Proposition 2.7, the type ${\rm II_{1}}$ factor $\operatorname{L}(\mathscr{R}_{i})$ is necessarily full by [Ch81].
$(\rm i)$ Assume that $\mathscr{R}=\mathscr{R}_{1}\times\cdots\times\mathscr{R}_{n}\cong\mathscr{T}_{1}\times\mathscr{T}_{2}$, where $\mathscr{T}_{1}$ and $\mathscr{T}_{2}$ are type ${\rm II_{1}}$ ergodic equivalence relations acting on $(Y_{1},\eta_{1})$ and $(Y_{2},\eta_{2})$ respectively. We may literally repeat the proof of Corollary B$(\rm i)$. There exists a partition $\{1,\dots,n\}=T_{1}\sqcup T_{2}$ into nonempty subsets such that $\operatorname{L}(\mathscr{T}_{1})\mathbin{\overline{\otimes}}\operatorname{L}^{\infty}(Y_{2})\preceq^{s}_{\operatorname{L}(\mathscr{R})}\operatorname{L}(\mathscr{R}_{T_{1}})\mathbin{\overline{\otimes}}\operatorname{L}^{\infty}(X_{T_{2}})$ and $\operatorname{L}^{\infty}(Y_{1})\mathbin{\overline{\otimes}}\operatorname{L}(\mathscr{T}_{2})\preceq^{s}_{\operatorname{L}(\mathscr{R})}\operatorname{L}^{\infty}(X_{T_{1}})\mathbin{\overline{\otimes}}\operatorname{L}(\mathscr{R}_{T_{2}})$. A combination of [DHI16, Lemma 2.6(3)] and [OP07, Proposition 2.4(3)] implies that $\operatorname{L}(\mathscr{T}_{j})$ is amenable relative to $\operatorname{L}(\mathscr{R}_{T_{j}})$ inside $\operatorname{L}(\mathscr{R})$ for every $j\in\{1,2\}$. Note that $\operatorname{L}(\mathscr{R})$ is full by [Co75, Corollary 2.3] and so are $\operatorname{L}(\mathscr{T}_{1})$, $\operatorname{L}(\mathscr{T}_{2})$, $\operatorname{L}(\mathscr{R}_{T_{1}})$, $\operatorname{L}(\mathscr{R}_{T_{2}})$. Then [IM19, Lemma 5.2] implies that $\operatorname{L}(\mathscr{T}_{j})\preceq_{\operatorname{L}(\mathscr{R})}\operatorname{L}(\mathscr{R}_{T_{j}})$ for every $j\in\{1,2\}$. Reasoning as in the proof of [Ho15, Corollary E(1)], we infer that there exists $t>0$ such that $\mathscr{T}_{1}^{t}\cong\mathscr{R}_{T_{1}}$ and $\mathscr{T}_{2}^{1/t}\cong\mathscr{R}_{T_{2}}$.
$(\rm ii)$ The proof follows by combining the ones of item $(\rm i)$, Corollary B$(\rm ii)$ and [Ho15, Corollary E]. This concludes the proof of Corollary C.
3.5. Proof of Theorem D
We prove the following key intermediate result which is analogous to [Dr19, Proposition 4.1].
Proposition 3.8.
Let $n\geq 1$. For every $1\leq i\leq n$, let $k_{i}$ be a local field of characteristic zero, $\mathbf{G}_{i}$ a $k_{i}$-isotropic $k_{i}$-simple algebraic $k_{i}$-group and $\Gamma_{i}<\mathbf{G}(k)$ a Zariski dense discrete subgroup with Kazhdan’s property (T). Set $\Gamma=\Gamma_{1}\times\cdots\times\Gamma_{n}$. Let $\Gamma\curvearrowright(X,\nu)$ be an essentially free ergodic pmp action and set $\mathscr{R}=\mathscr{R}(\Gamma\curvearrowright X)$. Assume that $\mathscr{R}\cong\mathscr{T}_{1}\times\mathscr{T}_{2}$, where $\mathscr{T}_{1},\mathscr{T}_{2}$ are type ${\rm II_{1}}$ ergodic equivalence relations. Set $M=\operatorname{L}(\mathscr{R})$ and for every $j\in\{1,2\}$, set $P_{j}=\operatorname{L}(\mathscr{T}_{j})$.
Then there exist a partition $\{1,\dots,n\}=T_{1}\sqcup T_{2}$ into nonempty subsets such that $P_{j}\preceq^{s}_{M}\operatorname{L}(\Gamma_{T_{j}}\curvearrowright X)$ for every $j\in\{1,2\}$.
Proof.
For every $j\in\{1,2\}$, denote by $T_{j}\subset\{1,\dots,n\}$ a minimal subset for which $P_{j}\preceq^{s}_{M}M_{T_{j}}$, where $M_{T_{j}}=\operatorname{L}(\Gamma_{T_{j}}\curvearrowright X)$ and $\Gamma_{T_{j}}=\prod_{i\in T_{j}}\Gamma_{i}$. Since $P_{j}$ is a type ${\rm II_{1}}$ factor, we have $T_{j}\neq\emptyset$. Set $T=T_{1}\cup T_{2}$, $\Gamma_{T}=\prod_{i\in T}\Gamma_{i}$ and $M_{T}=\operatorname{L}(\Gamma_{T}\curvearrowright X)$. Arguing as in the proof of Corollary B, we deduce that $T=\{1,\dots,n\}$.
It remains to prove that $T_{1}\cap T_{2}=\emptyset$. By contradiction, assume that $T_{1}\cap T_{2}\neq\emptyset$ and let $i\in T_{1}\cap T_{2}$. Set $\Lambda=\prod_{i\neq j}\Gamma_{i}$ and $\mathscr{S}=\mathscr{R}(\Lambda\curvearrowright X)$. Arguing as in the proof of Corollary B, by minimality of $T_{1}$ and $T_{2}$ and [DHI16, Lemma 2.8(2)], we have that $P_{j}\npreceq_{M}^{s}\operatorname{L}(\Lambda\curvearrowright X)$ for every $j\in\{1,2\}$. By Fact 2.2$(\rm i)$, this further implies that $\mathscr{T}_{j}\npreceq_{\mathscr{R}}\mathscr{S}$ as subequivalence relations. This however contradicts Theorem 3.6. Therefore, we have $\{1,\dots,n\}=T_{1}\sqcup T_{2}$. ∎
Proof of Theorem D.
By combining Proposition 3.8 and [Dr19, Theorem 3.2], we infer that there exist a partition $\{1,\dots,n\}=T_{1}\sqcup T_{2}$ into nonempty subsets, $t>0$, an essentially free ergodic pmp action $\Gamma_{T_{j}}\curvearrowright(X_{j},\nu_{j})$, where $\Gamma_{T_{j}}=\prod_{i\in T_{j}}\Gamma_{i}$ and $\operatorname{L}^{\infty}(X_{j},\nu_{j})=\operatorname{L}^{\infty}(X,\nu)^{\Gamma_{T_{j+1}}}$ for every $j\in\mathbb{Z}/2\mathbb{Z}$, a decomposition $\operatorname{L}(\mathscr{R})=\operatorname{L}(\mathscr{T}_{1})^{t}\mathbin{\overline{\otimes}}\operatorname{L}(\mathscr{T}_{2})^{1/t}$ and $u\in\mathscr{U}(\operatorname{L}(\mathscr{R}))$, such that $\Gamma\curvearrowright(X,\nu)$ is isomorphic to $\Gamma_{T_{1}}\times\Gamma_{T_{2}}\curvearrowright(X_{1}\times X_{2},\nu_{1}\otimes\nu_{2})$ and
$\operatorname{L}(\mathscr{T}_{1})^{t}=u\operatorname{L}(\Gamma_{T_{1}}\curvearrowright X_{1})u^{*}\quad\text{and}\quad\operatorname{L}(\mathscr{T}_{2})^{1/t}=u\operatorname{L}(\Gamma_{T_{2}}\curvearrowright X_{2})u^{*}.$
Arguing as in the proof of [Ho15, Corollary E], we infer that
$\mathscr{T}_{1}^{t}\cong\mathscr{R}(\Gamma_{T_{1}}\curvearrowright X_{1})\quad\text{and}\quad\mathscr{T}_{2}^{1/t}\cong\mathscr{R}(\Gamma_{T_{2}}\curvearrowright X_{2}).$
This finishes the proof of Theorem D. ∎
3.6. Proof of Corollary E
By applying Theorem D finitely many times, we can find a partition $\{1,\dots,n\}=T_{1}\sqcup\cdots\sqcup T_{r}$ into nonempty subsets, an essentially free ergodic pmp action $\Gamma_{T_{j}}\curvearrowright(X_{j},\nu_{j})$, where $\Gamma_{T_{j}}=\prod_{i\in T_{j}}\Gamma_{i}$ for every $j\in\{1,\dots,r\}$, such that $\Gamma\curvearrowright(X,\nu)$ is isomorphic to $\Gamma_{T_{1}}\times\cdots\times\Gamma_{T_{r}}\curvearrowright(X_{1}\times\cdots\times X_{r},\nu_{1}\otimes\cdots\otimes\nu_{r})$ and $\mathscr{R}(\Gamma_{T_{j}}\curvearrowright X_{j})$ is prime for every $j\in\{1,\dots,r\}$.
To prove the uniqueness part, we consider another partition $\{1,\dots,n\}=S_{1}\sqcup\cdots\sqcup S_{p}$ into nonempty subsets, an essentially free ergodic pmp action $\Gamma_{S_{j}}\curvearrowright(Y_{j},\nu_{j})$, where $\Gamma_{S_{j}}=\prod_{i\in S_{j}}\Gamma_{i}$ for every $j\in\{1,\dots,p\}$ such that $\Gamma\curvearrowright(X,\nu)$ is isomorphic to $\Gamma_{S_{1}}\times\cdots\times\Gamma_{S_{p}}\curvearrowright(Y_{1}\times\cdots\times Y_{p},\eta_{1}\otimes\cdots\otimes\eta_{p})$ and $\mathscr{R}(\Gamma_{S_{j}}\curvearrowright Y_{j})$ is prime for every $j\in\{1,\dots,p\}$. Without loss of generality, we may assume that $r\geq p$. We claim that for any $1\leq j\leq r$, there is $1\leq\ell\leq p$ such that $T_{j}\subset S_{\ell}$. To see this, note that for any $1\leq j\leq r$, there is $1\leq\ell\leq p$ such that $S_{\ell}\cap T_{j}\neq\emptyset$. Note that the action $\Gamma_{T_{j}}\curvearrowright(X,\nu)$ is isomorphic to $\Gamma_{S_{1}\cap T_{j}}\times\cdots\times\Gamma_{S_{p}\cap T_{j}}\curvearrowright(Y_{1}\times\cdots\times Y_{p},\eta_{1}\otimes\cdots\otimes\eta_{p})$. If $T_{j}\not\subset S_{\ell}$, it is easy to see that $\mathscr{R}(\Gamma_{T_{j}}\curvearrowright X_{j})$ is not prime, contradiction.
The above claim implies that $r=p$, and hence, $T_{j}=S_{j}$, for any $1\leq j\leq r$. Then the conclusion follows.